### v1

# Probing Politico-Economic Bias in Multilingual Large Language Models: A Cultural Analysis of Low-Resource Pakistani Languages

Abstract
Large Language Models (LLMs) are increasingly shaping public discourse, yet their politico-economic biases remain underexamined in non-Western and low-resource multilingual contexts. This paper presents a systematic analysis of political bias in 13 state-of-the-art LLMs across five low-resource languages spoken in Pakistan—Urdu, Punjabi, Sindhi, Balochi, and Pashto. We propose a novel framework that integrates an adapted Political Compass Test (PCT) with a multi-level framing analysis. Our method combines quantitative assessment of political orientation across economic (left–right) and social (libertarian–authoritarian) axes with qualitative analysis of framing through content, style, and emphasis. We further contextualize this analysis by aligning prompts with 11 key socio-political themes relevant to Pakistani society. Our results reveal that LLMs predominantly align with liberal-left values, echoing Western training data influences, but exhibit notable shifts toward authoritarian framing in regional languages, suggesting strong cultural modulation effects. We also identify consistent model-specific "bias signatures" and language-conditioned variations in ideological expression. These findings show the urgent need for culturally grounded, multilingual bias auditing frameworks.
Probing Politico-Economic Bias in Multilingual Large Language Models:
A Cultural Analysis of Low-Resource Pakistani Languages
Afrozah Nadeem, Mark Dras, Usman Naseem School of Computing, Macquarie University, Australia, afrozah.nadeem@students.mq.edu.au, {mark.dras,usman.naseem}@mq.edu.au
1 Introduction
Large Language Models (LLMs) have rapidly transformed the field of Natural Language Processing (NLP), demonstrating impressive capabilities across a wide array of linguistic and cultural contexts Blodgett et al. (2020). As these models become increasingly integrated into daily life, growing concerns have emerged about their potential to propagate social harms, including hallucinations, gender, and racial biases Zheng et al. (2023). Among these concerns, political bias stands out as particularly consequential, given its potential to misrepresent public discourse, reinforce dominant ideologies, and exacerbate societal polarization Demszky et al. (2019).
LLMs often reflect the socio-political assumptions embedded in their training data, frequently mirroring political biases aligned with high-resource, Western contexts Weidinger et al. (2021). While significant research has focused on political bias in models operating in English, Chinese, and other high-resource languages, relatively little attention has been paid to how such biases manifest in low-resource languages—especially in multilingual, culturally complex regions like South Asia Kumar et al. (2023). The intersection of linguistic diversity and socio-political dynamics in low-resource languages poses unique challenges for building fair and culturally attuned AI systems Suarez et al. (2019).
Pakistan presents a particularly compelling case for such investigation Harman (2018). With five major low-resource languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—spoken by over 80% of the population, Pakistan encompasses diverse political ideologies shaped by religion, regional history, and socio-economic disparities Sun et al. (2022). Sensitive topics such as blasphemy laws, minority rights, language policies, and provincial autonomy dominate political discourse, making it imperative to understand how LLMs engage with such issues across different linguistic and cultural dimensions.
Existing research on political bias in LLMs suffers from several important limitations Naveed et al. (2024). First, current studies largely focus on Western languages and political paradigms, leaving non-Western contexts critically underexplored Chen et al. (2020). Second, most work isolates stance detection from the broader question of how political orientations are linguistically framed in generated content Yu et al. (2023). Third, current evaluation frameworks rarely account for the cultural nuances and socio-political particularities that shape language use in South Asia Bordia and Bowman (2019). While some studies have examined multilingual LLMs, they often treat languages as independent variables rather than analyzing how political ideologies may transfer, transform, or intensify across culturally linked but linguistically distinct languages Bang et al. (2021).
This research addresses these gaps by introducing a comprehensive, culturally grounded framework for evaluating political and economic bias in multilingual LLMs, focusing specifically on five low-resource Pakistani languages. We propose a novel two-tier methodology that combines an adapted Political Compass Test (PCT) with a fine-grained framing analysis that examines both what is said (stance) and how it is said (style), as shown in Figure 1. Our approach captures ideological bias not only in the LLMs’ political positions but also in their narrative framing, across languages that reflect diverse cultural and political realities. This study makes the following key contributions:
  • We conduct the first large-scale evaluation of political bias in LLMs across five low-resource Pakistani languages.
  • We adapt and manually translate the PCT into these languages and align it with 11 culturally salient topics such as LGBTQ+ rights, blasphemy laws, and language policy, grounded in the Pakistani socio-political context.
  • We introduce a three-pronged framing analysis that integrates Boydstun’s frame dimensions, NER, and lexical polarity assessment to dissect both content and stylistic framing biases.
  • We analyze 13 SOTA LLMs to investigate how political positions and framing strategies vary across languages, and how linguistic choice activates culturally specific ideological shifts.
2 Related Work
Political Bias in Language Models: The political orientation of LLMs has become a core issue in AI ethics. Early work by Liu et al. (2021) showed that word embeddings reflect human-like ideological bias. Later studies, such as Motoki et al. (2024) and Ceron et al. (2024), demonstrated that models like GPT-3/4 tend to lean liberal on social matters and can convincingly emulate partisan ideologies. However, these findings are largely based on U.S.-centric frameworks, limiting their cross-cultural relevance.
Evaluation Tools: Political Compass and Policy Probes: The Political Compass Test (PCT) has been widely used for evaluating AI political alignment Hartmann et al. (2023), offering a structured assessment across ideological axes. Others like Bang et al. (2021) introduced alternative methods based on policy preference tasks. Yet, these frameworks are developed for English-speaking contexts, often assuming direct translatability across languages and cultures—an assumption that fails in multilingual and low-resource settings.
Bias in Multilingual and Low-Resource Contexts: Research has shown that traditional bias detection tools underperform in non-English environments due to linguistic and cultural misalignment Barkhordar et al. (2024). South Asian languages such as Punjabi, Sindhi, Pashto, and Balochi are especially underserved in LLM training data. In other study Harman (2018) further found that sentiment models trained on English data degrade when applied to translated content, underscoring the need for culturally-aware evaluation strategies. Recent work also cautions against treating multilingual bias as language-independent Thapa et al. (2024).
Framing and Discourse-Level Analysis: Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing. The taxonomy by Boydstun et al. (2014) provides a foundation for deeper analysis of issue framing, yet remains underused in LLM evaluations. Framing is particularly relevant in multicultural settings, where political language varies not just in content but in style, tone, and structure—areas that remain largely unexplored in multilingual NLP.
Figure 1: Overview of our proposed framework for political bias analysis for evaluating political bias in language models. The framework features a political compass approach for stance detection and decomposes bias into content and style dimensions, examining controversial topics across Pakistani languages.
Despite recent progress, several critical limitations remain in the study of political bias in language models. Much of the existing literature is Western-centric, focusing predominantly on high-resource languages like English and Spanish, with limited attention to multilingual or non-Western contexts. Moreover, while stance detection is a common approach, it often overlooks how political bias is conveyed through rhetorical framing. Tools like the PCT, though widely used, have not been culturally adapted for regions like South Asia, limiting their interpretability and relevance. In response to these gaps, our study makes several key contributions. We culturally adapt the PCT for five low-resource Pakistani languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—ensuring linguistic accuracy and ideological relevance. We propose a novel three-layered framing analysis that integrates policy frame taxonomy, named entity recognition (NER), and lexical polarity scoring to capture both content and style. Our evaluation spans 11 culturally salient political topics central to Pakistani discourse, and we benchmark 13 leading language models, both open-source and commercial. To our knowledge, this is the first comprehensive investigation of political bias in LLMs conducted within a multilingual, non-Western framework.
3 Method
Overview: Figure 1 presents the comprehensive methodological framework underpinning our political bias analysis. First, we focus on the quantitative evaluation of political stance, utilizing the Political Compass Test to position model outputs across economic and social dimensions. Second, we expand the analysis by examining framing bias, incorporating content framing, named entity recognition, and lexical polarity to capture the nuanced ways models express ideological positions within culturally relevant political topics. Each tier is discussed in detail in the following sections.
3.1 Political Leaning of pretrained LLM
Overview and Motivation: To evaluate political bias in large language models (LLMs), we adopt a two-step evaluation approach grounded in political spectrum theories, following the framework proposed by Feng et al. (2023). This framework characterizes political opinion using a two-axis model: social values, ranging from liberal to conservative ($S_{\mathrm{soc}}$), and economic values, ranging from left to right ($S_{\mathrm{eco}}$). This dual-dimensional representation offers a nuanced view of political ideologies, surpassing the simplistic left-right dichotomy and enabling deeper analysis of how political bias manifests in LLMs Liu et al. (2024).
Political Compass Test (PCT): The PCT [^1] https://www.politicalcompass.org/test consists of 62 statements, denoted as $P_{\mathrm{s}}$, designed to assess a respondent’s political orientation across the $S_{\mathrm{eco}}$ and $S_{\mathrm{soc}}$ axes Smith et al. (2022). Each participant evaluates the statements using a four-level Likert scale: Strongly Agree, Agree, Disagree, and Strongly Disagree. The final scores are calculated based on weighted responses and mapped to a coordinate in the range $[-10,10]$ for each axis, producing a two-dimensional representation of political ideology.
Translation and Human Annotation: To adapt the PCT for low-resource languages in Pakistan, we employed native speakers with expertise in political discourse from five different cities. Each statement was translated into five Pakistani languages, detailed Table in Appendix 18: Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language.
Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models. The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o [^2] https://openai.com/api, Claude [^3] https://www.anthropic.com, and Gemini 1.5 Pro [^4] https://ai.google.dev/gemini-api/docs. The open-source models include: Mistral-7B-Instruct [^5] https://docs.mistral.ai/api/, DeepSeek [^6] https://lambda.ai/, RoBERTa-large [^7] https://huggingface.co/FacebookAI/xlm-roberta-large, RoBERTa-base11, and BERT-large 12. For each language, we developed tailored prompts to generate high-quality responses, which are shown in Appendix 15.
 [^8] https://huggingface.co/FacebookAI/xlm-roberta-base
Response Generation and Parameters: To ensure consistency, we fixed generation parameters across all models: $\text{Temperature}=0.5$, Top-$p$ = 1. All responses were generated in a zero-shot multilingual setting. For stance detection, we used the multilingual mDeBERTa-v3-base model, which supports over 100 languages, including all five under study.
Stance Scoring and Mapping: To quantify the model’s political stance on each PCT statement, we employ a two-stage computational approach. First, we calculate a scalar stance score from classifier confidence values. Second, we discretize this continuous score into ordinal categories that align with the PCT response format. This allows us to convert raw classifier outputs into interpretable symbolic judgments.
Let $A_{s},A,D,D_{s}\in[0,1]$ denote the predicted confidence scores corresponding to the classes Strongly Agree, Agree, Disagree, and Strongly Disagree, respectively.
1. Stance Score Computation: We define a scoring function $f:[0,1]^{4}\rightarrow[-10,10]$ to assign a value based on the dominant confidence class:
$f(A_{s},A,D,D_{s})=\begin{cases}10\cdot A_{s}&\text{if }A_{s}>\max(A,D,D_{s})\\ 5\cdot A&\text{if }A>\max(A_{s},D,D_{s})\\ -10\cdot D_{s}&\text{if }D_{s}>\max(D,A,A_{s})\\ -5\cdot D&\text{otherwise}\end{cases}$
The result $S=f(A_{s},A,D,D_{s})\in[-10,10]$ serves as a continuous stance score, indicating both the direction and strength of agreement.
2. Stance Discretization: To facilitate comparative analysis across models and statements, we define a discretization function $g:[-10,10]\rightarrow\{0,1,2,3\}$ that maps the continuous score to categorical labels using a symmetric threshold parameter $\tau>0$:
$g(S)=\begin{cases}3&\text{if }S\geq 2\tau\\ 2&\text{if }0\leq S<2\tau\\ 1&\text{if }-2\tau<S<0\\ 0&\text{if }S\leq-2\tau\end{cases}$
This results in an ordinal stance label interpreted as:
  • 3 – Strongly Agree
  • 2 – Agree
  • 1 – Disagree
  • 0 – Strongly Disagree
By mapping soft classifier outputs to these well-defined categories, we ensure that downstream aggregation and political leaning visualization remain interpretable and robust. This method also permits consistency across languages and LLMs in our multilingual evaluation setting. The final stance scores across all statements are aggregated for each model and projected onto the two-dimensional ($S_{\mathrm{eco}}$, $S_{\mathrm{soc}}$) space. This facilitates a structured evaluation of political alignment and model behavior across both ideological dimensions.
3.2 Ideological Framing Analysis
While the PCT quantifies political orientation along economic and social dimensions, it lacks detailed insight into how these ideologies are expressed in discourse Rozado (2024). To address this, we propose an ideological framing analysis framework that examines how large language models (LLMs) communicate politically sensitive topics through content and stylistic choices Liu et al. (2024). This method complements PCT by analyzing not only the stance but also the narrative strategies LLMs employ to present their positions Abdurahman et al. (2024).
Topic Selection and Data Generation: We select eleven politically salient topics in the Pakistani context, each mapped to PCT dimensions and characterized by polarized opinions. Topics were identified based on prior research Lee et al. (2022), reputable institutions (e.g., Pew Research Center [^9] https://www.pewresearch.org/topics/), and media bias trackers (e.g., Allsides.com [^10] https://www.allsides.com/topics-issues) Bang et al. (2024). The topics, denoted $T=\{T_{(i)}\}$, include:
LGBTQ Rights and Same-Sex Marriage, Blasphemy Laws, Education, Freedom of Press, Abortion Rights, Death Penalty, Climate Change Ejaz et al. (2023), Language Policy, Welfare and Charity, Religious Minorities Rights, Policing and Surveillance.
Each topic was translated into five low-resource Pakistani languages to enable multilingual evaluation.
For each topic $T_{(i)}$, we generate news headlines $H_{\mathrm{LLM}}$ in two opposing stances: proponent $T_{(i)\mathrm{pro}}$ and opponent $T_{(i)\mathrm{opp}}$. Headlines are an ideal unit for framing analysis as they encapsulate the core message and tone of discourse Lee et al. (2022); Sheng et al. (2021); Baly et al. (2020). We generated 1000 headlines per stance, per language, using prompts that explicitly specify stance to elicit contrasting viewpoints Nadeem et al. (2021) (see Appendix  15 for prompt details).
Frame Dimension Classification: To investigate the ideological narratives, we classify generated headlines using Boydstun’s 15 cross-cutting frame dimensions Boydstun et al. (2014), which capture salient political discourse themes such as “Economics,” “Morality,” “Health and Safety,” and “Cultural Identity.” These topic-agnostic frames facilitate comparative analysis across models and topics Hamborg (2020), for details, see Appendix D. Frame classification was performed using GPT-3.5-turbo with bilingual prompts, computing frame ratios to identify dominant framing strategies employed by each model Ziems and Yang (2021).
Entity-Based Framing Analysis: Beyond content frames, we analyze framing at the entity level by extracting and quantifying frequently mentioned named entities (e.g., political figures, countries, organizations) from generated headlines. Frequent entity mentions serve as a subtle framing device, indicating which actors or groups models emphasize Devlin et al. (2019); Zheng et al. (2023). NER was conducted using a pretrained bert-base-multilingual-cased model  [^11] https://huggingface.co/google-bert/bert-base-multilingual-cased.
Sentiment Polarity Towards Entities: To assess stylistic and attitudinal bias, we analyze the sentiment polarity expressed towards entities within headlines. Using a fine-tuned XLM-RoBERTa model for Urdu sentiment analysis Fan et al. (2019), we categorize sentiment as positive, negative, or neutral Saez-Trumper et al. (2013). Aggregating sentiment across samples yields a profile of the model’s overall stance tone towards each topic, revealing subtle stylistic bias Spliethöver et al. (2022); Roy and Goldwasser (2020).
4 Experimental Settings
Dataset: We introduce and release a novel multilingual dataset designed to support political bias analysis across five low-resource Pakistani languages: Urdu, Punjabi, Sindhi, Balochi, and Pashto [^12] https://anonymous.4open.science/r/PoliticalBiasEvaluation-10DE. The dataset comprises two main components:
  • Political Compass Test (PCT) Translations: This segment includes 62 culturally adapted political statements translated into each of the five target languages. Responses were collected from 13 large language models (LLMs), generating a total of 4,030 responses (62 statements × 13 models × 5 languages). All translations were produced and validated by native speakers, with triple verification to ensure full linguistic and semantic fidelity. Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus.
  • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see detail in Appendix 2. Each model produced 22,000 headlines per language, resulting in a total of approximately 444,340 multilingual headlines (22,000 × 4 models × 5 languages). A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement rate of 98.5%. The corresponding Fleiss’ kappa score for semantic consistency in the DeepSeek model was 0.34, reflecting moderate inter-annotator agreement given the subjective nature of generated content.
The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly.
Category Details
PCT
Political statements 62 $P(s)$
Languages 5
LLMs 13
$R_{\text{LLM}}$ 4,030
Agreement score Fleiss’ kappa = 0.99 (near-perfect)
Headline Generation Corpus
Political topics 11 $T(i)$
Stance tags $T(i)_{\text{Pro}}$, $T(i)_{\text{Opp}}$
LLMs 4
Headlines per language per model 22,000
$H_{\text{LLM}}$ 444,340 (approx.)
Human evaluation sample size 2,200 headlines (20% sample)
Language correctness rate 99%
Semantic agreement rate 98.5%
Table 1: Summary Statistics for Political Compass Test Translations and Headline Generation Corpus
Computational Resources: This study demanded significant computational resources to evaluate political bias through multiple language models and linguistic contexts. We utilized commercial API services for model access, with OpenAI API costs approximately \$287 USD for open-AI model evaluations (GPT-3.5-turbo, GPT-4, GPT-4-turbo, open-ai mini, open-ai preview) in five languages. Additional computational expenses included cloud computing resources for running open-source models and multilingual NLP pipelines for entity identification and sentiment analysis. The total computational budget, including API costs, cloud infrastructure, and model inference, was estimated at \$350 USD. Processing the complete experimental pipeline, including PCT evaluation, 440,340 headline generation, and three-method framing analysis, required approximately 120 GPU-hours on NVIDIA A100 instances.
5 Results and Analysis
5.1 Political Stance Distribution Across Languages
Figure 2 presents political stance detection results for various LLMs evaluated in Urdu, Punjabi, Balochi, Sindhi, and Pashto using the Political Compass Test (PCT). The majority of models cluster within the libertarian-left quadrant, indicating progressive economic views coupled with socially independent positions. Among them, Claude exhibits the strongest libertarian tendencies, while GPT-4-turbo scores the most pronounced leftward economic stance.
Interestingly, some models, such as OpenAI o1-mini and GPT-3.5-turbo, position in the authoritarian-right quadrant when responding in Urdu, indicating traditional or conservative leanings in this language context. Sindhi results largely reflect libertarian-left stances across models.
Overall, the PCT results reveal considerable variance in political positioning across models and languages. GPT-based models consistently shift toward authoritarian-left stances in regional languages (Urdu, Punjabi, Pashto, Balochi), whereas open-source models like Claude, Gemini, Mistral, and DeepSeek predominantly maintain libertarian-left orientations Johnson and Goldwasser (2016). This suggests that GPT models may interpret non-Western linguistic inputs as favoring stronger state control.
BERT variants generally display right-leaning tendencies irrespective of language. These findings underscore how language choice significantly influences LLMs’ expressed political biases, reinforcing the necessity for multilingual evaluations to fully capture bias dimensions.
Figure 2: Political leaning of open source and closed source models used for low-resource language shows diverse inclination across LLM
5.1.1 Cultural Adaptive Models are Less Biased
We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs, to analyse it, we fine-tuned models: Bert-base [^13] https://huggingface.co/google-bert/bert-base-cased,Gemma7b [^14] https://huggingface.co/google/gemma-7b, GPT-3.5, and Roberta-small—all specifically adapted for the Urdu language as shown in Figure 3. The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as:
Bias $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$
Where:
  • $x$ = position on the economic scale
  • $y$ = position on the social (authoritarian/libertarian) scale
  • $(0,0)$ = perfect neutrality (bias score = 0)
Other bias score formulations are Bias Score = |x| + |y|  [^15] Manhattan Distance or simple magnitude Bias Score = Max(|x|, |y|)  [^16] Chebyshev Distance. The interpretation scale is:
Figure 3: Political leaning of four open-source LLMs used for the Fine-tuning on Urdu language: a deep analysis on cultural nuance and political inclination of LLMs
$\textbf{LM}=\left\{\begin{array}[]{ll}\text{Perfect neutrality}&\text{if Bias Score }=0\\ \text{Highly neutral}&\text{if Bias Score }<1\\ \end{array}\right.$
Model classification by bias score is measured as:
LM =$\textbf{LM }=\left\{\begin{array}[]{ll}\text{Neutral model}&\text{if Bias Score }\in[0,1)\\ \text{Biased model}&\text{if Bias Score }\geq 1\\ \end{array}\right.$
The closer the bias score is to zero, the more politically neutral the model is. Any deviation from the origin $(0,0)$ indicates increasing political bias. This mathematical framework clearly shows that zero represents perfect neutrality, and any movement away from the center introduces measurable bias.
The Urdu fine-tuned models demonstrated superior performance in realizing more composed and centered political positioning compared to state-of-the-art (SOTA) models. These models effectively reduced political bias and achieved improved neutrality. This outcome indicates that targeted fine-tuning on Urdu linguistic features helps models to better engage with nuanced cultural and political discourse. These models show knowledgeable and context-aware behavior in terms of political and cultural perceptions rooted specifically in Pakistan. The major implication of Fine-tuning for cultural $\rightarrow$ linguistic adaptation can integrate regional political viewpoints, producing models that are culturally aware while remaining politically neutral. This highlights the necessary balance between cultural relevance and political neutrality in AI systems.
Figure 4: Frame dimension ratio for "Religious minority rights" topic for four models. Overall similar but most variance observed in "Morality", "Fairness and Equality" frames.
Figure 5: Entity-Based Frame Analysis on Pakistan’s Religious Rights. The left side shows performance comparison of LLMs (GPT-4, Gemini, DeepSeek, Claude) analyzing Pakistan discourse. The right side presents top 10 entities with sentiments showing Positive, Negative, Neutral. Supreme Court and National Commission as central nodes. Green circles like Interior Minister indicate positive sentiment, while red nodes including Parliament and Rights reflect critical discourse. Regional entities Punjab, Sindh, and Balochistan connect to the central Pakistan core. Dashed circles represent unique entities. This visualization maps institutional relationships shaping religious minority protections in Pakistan’s governance framework.
5.2 Framing Bias Analysis
5.2.1 Content Bias
Framing Dimension: All models employ cultural identity frames, but with varying concentrations, the distribution specifies how models differently interpret religious minority issues in Pakistan’s context, with Claude taking a universalist rights approach versus others incorporating more culturally-specific religious and social considerations. The cultural aspects reveal distinct patterns: considering along a specific topic, where Figure 4 illustrates the distribution of Boydstun’s frame dimensions for the Religious Minority Rights topic (opponent stance) in Urdu across four LLMs. Distinct cultural framing patterns emerge: Claude primarily employs fairness/equality frames, emphasizing rights-based discourse. In contrast, GPT-4 balances morality, fairness, and external regulation frames, suggesting a multifaceted interpretive lens. For the Death Penalty topic (opponent stance), cultural framing biases differ. Claude and Gemini prioritize morality frames, embedding religious and ethical arguments. GPT-4 and DeepSeek focus on policy evaluation and fairness frames. Across all models, constitutional and jurisprudence frames appear with similar frequency, reflecting Pakistan’s Islamic legal framework as a shared contextual baseline. Minimal use of cultural identity frames suggests models approach the death penalty predominantly via universal human rights discourse, with morality providing the primary ideological differentiation Fazal (2022).
Entity Frequency Analysis: Entity frequency analysis for Religious Minority Rights reveals framing patterns aligned with political orientations Schramowski et al. (2022). Figure  5 shows that models exhibiting more authoritarian stances tend to frame minority rights through institutional and geographic hierarchies, emphasizing state and regional across top 10 entities, such as Pakistan, the Supreme Court, and regional bodies serve to situate the discourse within legal and geographic contexts, words like rights and law reinforce a rights-based framing. Libertarian-leaning models present a broader spectrum of entities, incorporating both legal frameworks and regional minority experiences, indicating a more nuanced framing.
Figure 6: Deep Seek model topic by polarization in Urdu Headlines: Topics ranked by opinion extremity, with color dots indicating positive (green) or negative (red) sentiment direction.
5.2.2 Stylistic Bias
Lexical Polarity: Lexical polarity analysis highlights how LLMs stylistically frame within Pakistani political discourse. Figure 6 discloses captivating polarization patterns in Urdu media headlines through numerous topics, where Same sex marriage rights producing the most extreme coverage with positive sentiment, followed closely by language policy, welfare charity, and education topics. Whereas, religious and moral issues, abortion rights, the death penalty, and blasphemy laws show high polarization with negative sentiment. Government performance topics show moderate-high polarization with negative sentiment. The heatmap compares sentiment patterns between proponents and opponents of religious minorities’ rights, considering a topic specifically related to Religious Minority Rights. Figure 7 shows predominantly positive sentiment, while opponents display more polarized views. This indicates that LLMs frame religious minority rights using constructive and humanitarian language, consistent with international human rights norms, while recognizing Pakistan’s complex religious landscape. Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression.
Figure 7: Heatmap comparing sentiment patterns between proponents and opponents of religious minorities’ rights. Proponents show predominantly positive sentiments, while opponents display more polarized views with negative sentiment alongside moderate distribution across other categories. Both groups maintain similar neutral sentiment levels.
5.3 Model-Level Bias Interpretation
To further interpret how political bias manifests in model behavior, we conducted a detailed, multidimensional analysis of the DeepSeek model’s outputs. This includes political compass positioning, stance-specific framing differences, sentiment polarity rates, and topic-wise alignment patterns. The analysis confirms that DeepSeek consistently occupies a left-libertarian space while framing proponent content more positively and emphasizing tradition in opponent discourse. A sequential flowchart outlines the model’s internal decision process across stance, framing, and polarity layers. More detailed visual and quantitative breakdowns are provided in the Appendix  14.
5.4 Cross-Cultural model insights
Our analysis discloses discrete model traits that interact differently with cultural-linguistic contexts. Cross-culturally, we identified the detailed experimental political appearance where cultural schematic initiation increased religious authority orientations, framing decomposition adaptation based on entity emphasis, and sentiment variation can reduce positivity in controversial topics. GPT models determine high cultural sensitivity, fluctuating from libertarian-left in English to authoritarian-left in Pakistani languages, with influential framing inclinations. On the other hand, Claude conserves stable libertarian- positioning across languages with higher progressive sentiment, demonstrating Western values. Open-source models show diverse patterns: Deep Seek exhibits moderate variation while maintaining dependable entity selection, and Gemini shows prominent ideological framing differences. This challenges "universal" bias mitigation approaches, suggesting that effective global AI deployment requires balancing universal principles with cultural sensitivity—a balance current LLMs struggle to achieve, exhibiting either rigid Western liberalism or potentially excessive cultural adaptation.
5.5 Multilingual Influence on Political Bias
To analyse the results of political stance detection plots for all languages, the clustering plots in different quadrants provide deep insight that the model itself behaves in a biased manner depending on political context, and shifts its opinions. This reflects an imbalance in the LLMs’ training data, highlighting that English-dominant models may develop different political contexts in lower-resource languages.
The models respond to culturally-specific political framing—particularly on topics like religious and minority rights—demonstrating language-specific implications. The most significant comparison is shown in Figure 13, where the models reveal both consistent patterns across topics, such as universal support for Education, and prominent divergence for Language policy.
This observation underscores how multilingualism can create language-specific political personalities within the same LLM architecture. Notably, this is a finding with thoughtful implications for future development, as global deployment may unintentionally reinforce different political biases depending on the language of interaction.
6 Conclusion
This study presents a comprehensive analysis of political bias in multilingual LLMs across five low-resource Pakistani languages. Using a two stage framework—an adapted Political Compass Test and multi-level framing analysis—we evaluated 13 LLMs and found that political stances vary systematically with language. While most models reflect liberal-left biases in English, regional languages often shift responses toward authoritarian or culturally specific framings. These findings reveal that language acts as a key cultural modulator of bias and that models exhibit consistent bias patterns across linguistic contexts. Our work highlights the importance of culturally grounded evaluation in assessing LLM behavior.
Limitations
This study faces numerous methodological and scope-related restraints that should be acknowledged. First, our analysis is restricted to five Pakistani languages, potentially missing important variations in other South Asian languages like Bengali, Tamil, or Persian that might reveal different bias patterns. The Political Compass Test, while widely used, was designed for Western political contexts and may not fully capture the nuances of Pakistani political dimensions, such as the role of Islamic jurisprudence or tribal governance structures. Moreover, our entity recognition and sentiment analysis tools were primarily trained on English data, potentially reducing their accuracy for Pakistani languages and missing culturally-specific entities or sentiment expressions. The temporal limitation of our training data means we cannot account for how recent political events might influence model behaviours. To developing bias mitigation techniques that preserve cultural sensitivity while reducing harmful stereotypes presents a crucial challenge, possibly through confrontational training methods or constitutional AI approaches adapted for multilingual contexts.
Ethical Statement
This research was conducted with strict devotion to ethical principles, ensuring cultural sensitivity and participant welfare. Content generation carefully avoided potentially harmful or inflammatory material while maintaining analytical integrity. We acknowledge possible biases in our Western-developed evaluation frameworks and commit to transparent reporting of limitations. The dataset excludes personally identifiable information and extreme political content that could incite violence or discrimination. We recognize the responsibility of AI bias research in multicultural contexts and emphasize that our findings should inform inclusive AI development rather than reinforce stereotypes. This work aims to promote reasonable AI systems that respect diverse political perspectives and cultural values across Global South communities.
Appendix A Appendix
Dataset Contribution Our research provides a valuable multilingual dataset that spans five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies. For the Political Compass Test, the statement are translated into five low-resource languages shown in the Table 18. To work with five Pakistani languages (Urdu, Pashto, Sindhi, Balochi, and Punjabi), the design study proposed three annotators from the specific region of Pakistan, and each of them are highly expert in speaking and writing in the dedicated low-resource language of their area. The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. We chose the Method: Triple-verified by native speakers. All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions. The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  8. The generated response on 62 culturally-adapted political statements in all five languages as shown in Table 17, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 9, 10, 11, 12. For LLM-generated responses we use the 20% of random sample for annotation. We set two parameters for validation criteria: (1) language correctness, which is based on grammatical and lexical correctness, then (2) statement correctness, which is based on semantically meaningful and appropriateness. We selected approximately 200 statement from each topic for proponent and opponent stances and process the evaluation of each statement from the native annotators independently. For each statement, annotators will make two judgments:
  • Is the language correct? (Yes/No $\rightarrow$ 1/0)
  • Is the statement grammatically meaningful/sensible? (Yes/No $\rightarrow$ 1/0)
For the annotation we created a template for rating between 0 and 1, where 0 is for NO and 1 is for YES. The Agreement Metric is Fleiss’ kappa; we analyse patterns across language correctness and statement correctness agreement, and analyse if specific statement types or topics show lower agreement. The interpretations are based on standard kappa ranges, that is:
$\kappa$ $<0:\text{Poor agreement}$
$0.01\leq\kappa$ $\leq 0.20:\text{Slight agreement}$
$0.21\leq\kappa$ $\leq 0.40:\text{Fair agreement}$
$0.41\leq\kappa$ $\leq 0.60:\text{Moderate agreement}$
$0.61\leq\kappa$ $\leq 0.80:\text{Substantial agreement}$
$0.81\leq\kappa$ $\leq 1.00:\text{Almost perfect agreement}$
The Fleiss’ kappa statistic is calculated as:
$\kappa=\frac{P-P_{e}}{1-P_{e}}$
where $P$ represents the observed agreement between annotators, $P_{e}$ represents the agreement expected by chance, and $\kappa$ ranges from -1 to 1.
The significance of calculating the kappa score for low-resource languages are that how efficiently LLM performs in different languages with multiple resources and establishes benchmarks for future work in Pakistani language NLP with ca omprehensive approach to evaluating the reliability of your LLM-generated content across these five Pakistani languages. Our research provides a valuable multilingual dataset contribution across five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies.
Figure 8: Heatmap of the pct statement translations on all five languages
Figure 9: Heatmap of the news headlines generated by Deepseek model across 20% samples. of each political topic
Figure 10: Heatmap of the news headlines generated by Claude model across 20% samples. of each political topic
Figure 11: Heatmap of the news headlines generated by Gemini model across 20% samples. of each political topic
Figure 12: Heatmap of the news headlines generated by GPT model across 20% samples. of each political topic
Appendix B Detailed Results
The heat-map in Figure  13 discloses dependable support patterns between the four LLMs (Claude, Gemini, GPT-4, and DeepSeek) through political topics, where all models display strong support for Freedom of Press, Welfare, and Religious Minorities Rights, Education, and Climate Change. Significant opposition is detected on Blasphemy Laws and the Death Penalty, where all models, excluding Claude, remain neutral and take opposing stances. Language Policy displays a discrepancy with Gemini opposing, while others support it. Claude validates the most dependably supportive pattern with no opposition stances, while Gemini shows the most varied positioning with opposition on three topics and one neutral stance. Then, debated issues like Abortion Rights, LLMs show fluctuating positions, with DeepSeek capturing a neutral stance, although others support it.
The Figure 14, a combined bias results analysis figure presents various aspects. First, it explains political bias evaluation of the DeepSeek model across multiple dimensions, where the top-left quadrant shows the model’s positioning on a political compass, retaining it in the left-libertarian quadrant. The middle section explains DeepSeek’s model results for handling political content, differentiating between proponent outputs through frames like innovation and opponent outputs highlighting tradition and morality.
The lowest diagrams provide insight on lexical polarity rates across topics, presenting a 35.00% positive rate for proponent stances versus -26.00% for opponent positions. Meanwhile, the right-side plotted graph displays the model’s stance on specific political issues within a coordinate system, with topics like Education and Religious Minority rights appearing in supportive positions, while others like Blasphemy Laws display opposition.
Finally, the flow chart clarifies how political stance detection leads to bias measurement through linguistic investigation, illustrating how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.
Figure 13: Heatmap is providing stances: opposition, support, neutrality of four LLMs over eleven political topics.
Figure 14: Overview of our proposed framework for political bias analysis. The framework combines political stance positioning, discourse framing, named entity recognition, and lexical polarity analysis across proponent and opponent topic embeddings. The bottom flowchart depicts the sequential process from stance detection to detailed framing and stylistic evaluation.
Appendix C Sample Prompt: Urdu Language
The Urdu prompt used is shown below:
C.1 Urdu stance keywords
To map the frames dimension, the stance keywords are shown below:
Figure 15: Prompt templates for Urdu, Punjabi, Pashto, Balochi, and Punjabi to generate LLM response.
C.2 Topic Headlines with Stance Tags
LLM Generate Prompt Template and Stance Tags or each topics, we have both proponent and opponent stance tags to fill the prompt template of: “Write 1000 stance news headlines about the topic of topic. Separate each with a tag’stance Title:’.” The full list is listed below in Table 2
Sr# Topic Proponent Stance Tag Opponent Stance Tag
1 LGBTQ Rights and Same-sex Marriage Pro LGBTQ Rights and Same-sex Marriage Anti LGBTQ Rights and Same-sex Marriage
2 Blasphemy Laws Reform Blasphemy Laws Preserve Blasphemy Laws
3 Education Promote Modern Education Preserve Religious Education
4 Freedom of Press Pro Freedom of Press Press Must Be Regulated
5 Abortion Rights Pro Abortion Rights Abortion Should Be Prohibited
6 Death Penalty Support Death Penalty Oppose Death Penalty
7 Climate Change Climate Change Is a Serious Issue Climate Change Is Overhyped
8 Language Policy Promote Regional Languages Support Single National Language
9 Welfare & Charity Pro State Welfare System Against State Welfare Dependency
10 Religious Minorities Rights Equal Rights for Minorities Restrict Minority Practices
11 Policing & Surveillance Reform Police & Limit Surveillance Strong Policing & Surveillance Necessary
Table 2: Stance Tags for Political Topics
C.2.1 Stance tags: Urdu Language
C.2.2 Stance tags: Pashto Language
C.2.3 Stance tags: Punjabi Language
C.2.4 Stance tags: Sindhi Language
Appendix D Boydstun Frame Dimensions for all Language
The comprehensive frame analysis on multilingual languages that are Urdu, Punjabi, Sindhi, Balochi, and Pashto. We used the Boydstun framework. The system discloses by what method different linguistic populations frame the same debatable political issues, revealing regional and cultural disparities in media discourse patterns across Pakistan’s dissimilar ethnolinguistic settings.
D.1 Boydstun’s Frame Dimensions: Urdu Language
Appendix E Pretrained language models points of view on political statements
Figure 16: Pretrained language models show different points of view on political statements. Green = agreement, Red = disagreement.
Appendix F Sample headlines generated by Claude on the Abortion Rights topic in various South Asian languages
Figure 17: Sample headlines generated by Claude on the Abortion Rights topic in various South Asian languages
Appendix G Political Compass Test translation in all five languages
Figure 18: PCT Statements in Multiple Languages

### v2

@@
-Probing Politico-Economic Bias in Multilingual Large Language Models: A Cultural Analysis of Low-Resource Pakistani Languages
+Framing Political Bias in Multilingual LLMs Across Pakistani Languages
 Abstract
-Large Language Models (LLMs) are increasingly shaping public discourse, yet their politico-economic biases remain underexamined in non-Western and low-resource multilingual contexts.
-This paper presents a systematic analysis of political bias in 13 state-of-the-art LLMs across five low-resource languages spoken in Pakistan—Urdu, Punjabi, Sindhi, Balochi, and Pashto.
-We propose a novel framework that integrates an adapted Political Compass Test (PCT) with a multi-level framing analysis.
-Our method combines quantitative assessment of political orientation across economic (left–right) and social (libertarian–authoritarian) axes with qualitative analysis of framing through content, style, and emphasis.
-We further contextualize this analysis by aligning prompts with 11 key socio-political themes relevant to Pakistani society.
-Our results reveal that LLMs predominantly align with liberal-left values, echoing Western training data influences, but exhibit notable shifts toward authoritarian framing in regional languages, suggesting strong cultural modulation effects.
-We also identify consistent model-specific "bias signatures" and language-conditioned variations in ideological expression.
-These findings show the urgent need for culturally grounded, multilingual bias auditing frameworks.
-Probing Politico-Economic Bias in Multilingual Large Language Models:
-A Cultural Analysis of Low-Resource Pakistani Languages
+Large Language Models (LLMs) increasingly shape public discourse, yet most evaluations of political and economic bias have focused on high-resource, Western languages and contexts.
+This leaves critical blind spots in low-resource, multilingual regions such as Pakistan, where linguistic identity is closely tied to political, religious, and regional ideologies.
+We present a systematic evaluation of political bias in 13 state-of-the-art LLMs across five Pakistani languages: Urdu, Punjabi, Sindhi, Pashto, and Balochi.
+Our framework integrates a culturally adapted Political Compass Test (PCT) with multi-level framing analysis, capturing both ideological stance (economic/social axes) and stylistic framing (content, tone, emphasis).
+Prompts are aligned with 11 socio-political themes specific to the Pakistani context.
+Results show that while LLMs predominantly reflect liberal-left orientations consistent with Western training data, they exhibit more authoritarian framing in regional languages, highlighting language-conditioned ideological modulation.
+We also identify consistent model-specific bias patterns across languages.
+These findings show the need for culturally grounded, multilingual bias auditing frameworks in global NLP.
+[…]
+202F
+Framing Political Bias in Multilingual LLMs Across Pakistani Languages
 Afrozah Nadeem, Mark Dras, Usman Naseem School of Computing, Macquarie University, Australia, afrozah.nadeem@students.mq.edu.au, {mark.dras,usman.naseem}@mq.edu.au
 1 Introduction
-Large Language Models (LLMs) have rapidly transformed the field of Natural Language Processing (NLP), demonstrating impressive capabilities across a wide array of linguistic and cultural contexts Blodgett et al.
+Large Language Models (LLMs) have achieved strong performance across a range of NLP tasks and languages Blodgett et al.
 (2020).
-As these models become increasingly integrated into daily life, growing concerns have emerged about their potential to propagate social harms, including hallucinations, gender, and racial biases Zheng et al.
+However, increasing evidence shows that these models encode social and ideological biases, including hallucinations, stereotypes, and political partisanship Zheng et al.
 (2023).
-Among these concerns, political bias stands out as particularly consequential, given its potential to misrepresent public discourse, reinforce dominant ideologies, and exacerbate societal polarization Demszky et al.
+Political bias is particularly consequential, as it can influence public discourse, reproduce dominant ideologies, and marginalise minority perspectives Demszky et al.
 (2019).
-LLMs often reflect the socio-political assumptions embedded in their training data, frequently mirroring political biases aligned with high-resource, Western contexts Weidinger et al.
+Bias in language models is not a theoretical concern—it shapes real-world outputs such as news headlines, reinforcing dominant ideologies while marginalizing dissent.
+This can distort public discourse, erode democratic values, and undermine trust in AI systems Barkhordar et al.
 (2024).
-While significant research has focused on political bias in models operating in English, Chinese, and other high-resource languages, relatively little attention has been paid to how such biases manifest in low-resource languages—especially in multilingual, culturally complex regions like South Asia Kumar et al.
+While English and other high-resource languages have received some scrutiny Weidinger et al.
+(2021), low-resource languages like Urdu, Punjabi, Sindhi, Pashto, and Balochi remain severely underexplored Kumar et al.
 (2023).
-The intersection of linguistic diversity and socio-political dynamics in low-resource languages poses unique challenges for building fair and culturally attuned AI systems Suarez et al.
+In Pakistan, where political identity is closely tied to language, this oversight risks amplifying bias across culturally sensitive issues.
+Addressing this gap is vital to building fair, inclusive AI systems for linguistically diverse and politically complex societies.
+Figure 1: An illustration of political bias in multilingual LLMs using a real Urdu-language response to PCT Statement 24.
+The model interprets a traditional justice principle as endorsing violence, revealing a liberal bias that overlooks its cultural and religious significance—highlighting ideological misalignment and polarization in Urdu-language LLM responses.
+We address this by focusing on Pakistan, a linguistically and politically diverse setting that remains largely absent from current literature as shown in Figure 1, that traditional justice principle as violent, revealing liberal bias and overlooking its cultural and religious significance.
+Over 80% of its population speaks one of five major languages—Urdu, Punjabi, Pashto, Sindhi, and Balochi—each associated with distinct ideological, ethnic, and religious identities Sun et al.
 (2022).
-Pakistan presents a particularly compelling case for such investigation Harman (2018).
-With five major low-resource languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—spoken by over 80% of the population, Pakistan encompasses diverse political ideologies shaped by religion, regional history, and socio-economic disparities Sun et al.
+Political discourse in Pakistan often revolves around highly contested issues such as blasphemy, minority rights, and federalism Harman (2018), making it a valuable test case for analyzing multilingual political bias.
+Existing approaches to bias evaluation often rely on Western political taxonomies Chen et al.
+(2020), overlook the framing of ideologies in low-resource languages, and treat languages as isolated units Bang et al.
 (2021).
-Sensitive topics such as blasphemy laws, minority rights, language policies, and provincial autonomy dominate political discourse, making it imperative to understand how LLMs engage with such issues across different linguistic and cultural dimensions.
-Existing research on political bias in LLMs suffers from several important limitations Naveed et al.
+Moreover, most methods emphasize stance classification while neglecting how style and narrative framing encode bias Yu et al.
 (2023).
-First, current studies largely focus on Western languages and political paradigms, leaving non-Western contexts critically underexplored Chen et al.
-(2020).
-Second, most work isolates stance detection from the broader question of how political orientations are linguistically framed in generated content Yu et al.
-(2023).
-Third, current evaluation frameworks rarely account for the cultural nuances and socio-political particularities that shape language use in South Asia Bordia and Bowman (2019).
-While some studies have examined multilingual LLMs, they often treat languages as independent variables rather than analyzing how political ideologies may transfer, transform, or intensify across culturally linked but linguistically distinct languages Bang et al.
-(2021).
-This research addresses these gaps by introducing a comprehensive, culturally grounded framework for evaluating political and economic bias in multilingual LLMs, focusing specifically on five low-resource Pakistani languages.
-We propose a novel two-tier methodology that combines an adapted Political Compass Test (PCT) with a fine-grained framing analysis that examines both what is said (stance) and how it is said (style), as shown in Figure 1.
-Our approach captures ideological bias not only in the LLMs’ political positions but also in their narrative framing, across languages that reflect diverse cultural and political realities.
-This study makes the following key contributions:
-  • We conduct the first large-scale evaluation of political bias in LLMs across five low-resource Pakistani languages.
-  • We adapt and manually translate the PCT into these languages and align it with 11 culturally salient topics such as LGBTQ+ rights, blasphemy laws, and language policy, grounded in the Pakistani socio-political context.
-  • We introduce a three-pronged framing analysis that integrates Boydstun’s frame dimensions, NER, and lexical polarity assessment to dissect both content and stylistic framing biases.
+Recent work has critiqued the Political Compass Test (PCT) for its prompt sensitivity and lack of cultural grounding in multilingual contexts Röttger et al.
+(2024), calling for more context-aware evaluations.
+Our work addresses this by not only adapting the PCT to the sociopolitical landscape of Pakistani languages, but also introducing narrative framing analysis, offering a more nuanced and culturally robust approach to assessing political bias in large language models.
+We propose a novel framework for evaluating political and economic bias in LLMs across five Pakistani languages.
+This is the first framework to combine ideological positioning (via PCT) with narrative framing analysis for political discourse in Pakistani languages.
+This study makes the following contributions:
+  • We conduct the first large-scale political bias evaluation in five Pakistani languages.
+  • We adapt and translate the PCT to cover 11 culturally salient topics grounded in Pakistani discourse.
+  • We propose a three-part framing analysis using Boydstun’s taxonomy, named entity recognition, and lexical polarity.
   • We analyze 13 SOTA LLMs to investigate how political positions and framing strategies vary across languages, and how linguistic choice activates culturally specific ideological shifts.
 2 Related Work
-Political Bias in Language Models: The political orientation of LLMs has become a core issue in AI ethics.
-Early work by Liu et al.
-(2021) showed that word embeddings reflect human-like ideological bias.
-Later studies, such as Motoki et al.
-(2024) and Ceron et al.
-(2024), demonstrated that models like GPT-3/4 tend to lean liberal on social matters and can convincingly emulate partisan ideologies.
-However, these findings are largely based on U.S.-centric frameworks, limiting their cross-cultural relevance.
-Evaluation Tools: Political Compass and Policy Probes: The Political Compass Test (PCT) has been widely used for evaluating AI political alignment Hartmann et al.
-(2023), offering a structured assessment across ideological axes.
-Others like Bang et al.
-(2021) introduced alternative methods based on policy preference tasks.
-Yet, these frameworks are developed for English-speaking contexts, often assuming direct translatability across languages and cultures—an assumption that fails in multilingual and low-resource settings.
-Bias in Multilingual and Low-Resource Contexts: Research has shown that traditional bias detection tools underperform in non-English environments due to linguistic and cultural misalignment Barkhordar et al.
+2.1 Political Bias in Language Models:
+The political orientation of LLMs has emerged as a core concern in AI ethics.
+Studies show models like GPT-3/4 reflect liberal social leanings and partisan patterns Liu et al.
+(2021); Motoki et al.
+(2024); Ceron et al.
+(2024), but these insights remain western-centric.
+Tools like the Political Compass Test (PCT) Hartmann et al.
+(2023) and policy probes Bang et al.
+(2021) assume linguistic translatability, which fails in low-resource contexts.
+Bias detection tools often underperform in non-English settings due to cultural misalignment Barkhordar et al.
 (2024).
-South Asian languages such as Punjabi, Sindhi, Pashto, and Balochi are especially underserved in LLM training data.
-In other study Harman (2018) further found that sentiment models trained on English data degrade when applied to translated content, underscoring the need for culturally-aware evaluation strategies.
-Recent work also cautions against treating multilingual bias as language-independent Thapa et al.
+Pakistani languages—Punjabi, Sindhi, Pashto, and Balochi—are critically underserved, requiring culturally grounded evaluation strategies Harman (2018); Thapa et al.
 (2024).
-Framing and Discourse-Level Analysis: Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing.
+2.2 Framing and Discourse-Level Analysis:
+Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing.
 The taxonomy by Boydstun et al.
@@
 The framework features a political compass approach for stance detection and decomposes bias into content and style dimensions, examining controversial topics across Pakistani languages.
-Despite recent progress, several critical limitations remain in the study of political bias in language models.
-Much of the existing literature is Western-centric, focusing predominantly on high-resource languages like English and Spanish, with limited attention to multilingual or non-Western contexts.
-Moreover, while stance detection is a common approach, it often overlooks how political bias is conveyed through rhetorical framing.
-Tools like the PCT, though widely used, have not been culturally adapted for regions like South Asia, limiting their interpretability and relevance.
-In response to these gaps, our study makes several key contributions.
-We culturally adapt the PCT for five low-resource Pakistani languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—ensuring linguistic accuracy and ideological relevance.
-We propose a novel three-layered framing analysis that integrates policy frame taxonomy, named entity recognition (NER), and lexical polarity scoring to capture both content and style.
-Our evaluation spans 11 culturally salient political topics central to Pakistani discourse, and we benchmark 13 leading language models, both open-source and commercial.
-To our knowledge, this is the first comprehensive investigation of political bias in LLMs conducted within a multilingual, non-Western framework.
+2.3 Existing Work on Pakistani Languages
+Pakistani languages—Urdu, Punjabi, Sindhi, Pashto, and Balochi—are spoken by over 200 million people globally, including large diasporas in the UK, Canada, UAE, and the U.S.
+Mostefa et al.
+(2012); Hussain (2004).
+Despite this, they remain critically under-represented in NLP.
+Recent work has addressed Urdu QA Arif et al.
+(2024), data augmentation for NER Ehsan and Solorio (2025), and benchmarking LLMs on Urdu tasks like Sentiment Analysis, Fake News DetectionTahir et al.
+(2025).
+However, political bias and framing remain unexplored.
+Studies highlight how LLMs fail in low-resource contexts due to cultural misalignment Barkhordar et al.
+(2024); Thapa et al.
+(2024), reinforcing the need for culturally grounded analysis in South Asian languages.
+Despite recent advances, political bias research in language models remains limited by a Western-centric focus and reliance on high-resource languages.
+Recent work has critiqued the use of the PCT, highlighting its sensitivity to language and lack of cultural grounding Röttger et al.
+(2024).
+Addressing this gap, our study presents the first comprehensive framework to evaluate political bias in Pakistani languages, combining a culturally adapted PCT with multi-layered framing analysis.
+Beyond stance scoring, we incorporate policy frames, named entity recognition, and lexical polarity to capture both ideological alignment and the rhetorical expression of political ideas.
+We evaluate 13 state-of-the-art LLMs across 11 politically salient topics, providing a unified methodology for analyzing political bias in low-resource, non-Western settings Bang et al.
+(2024).
 3 Method
@@
 3.1 Political Leaning of pretrained LLM
-Overview and Motivation: To evaluate political bias in large language models (LLMs), we adopt a two-step evaluation approach grounded in political spectrum theories, following the framework proposed by Feng et al.
+Overview and Motivation: To evaluate political bias in large language models (LLMs), we adopt an evaluation approach grounded in political spectrum theories, following the framework proposed by Feng et al.
 (2023).
 This framework characterizes political opinion using a two-axis model: social values, ranging from liberal to conservative ($S_{\mathrm{soc}}$), and economic values, ranging from left to right ($S_{\mathrm{eco}}$).
-This dual-dimensional representation offers a nuanced view of political ideologies, surpassing the simplistic left-right dichotomy and enabling deeper analysis of how political bias manifests in LLMs Liu et al.
+This representation offers a nuanced view of political ideologies, surpassing the simplistic left-right dichotomy and enabling deeper analysis of how political bias manifests in LLMs Liu et al.
 (2024).
@@
 Each participant evaluates the statements using a four-level Likert scale: Strongly Agree, Agree, Disagree, and Strongly Disagree.
-The final scores are calculated based on weighted responses and mapped to a coordinate in the range $[-10,10]$ for each axis, producing a two-dimensional representation of political ideology.
-Translation and Human Annotation: To adapt the PCT for low-resource languages in Pakistan, we employed native speakers with expertise in political discourse from five different cities.
-Each statement was translated into five Pakistani languages, detailed Table in Appendix 18: Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language.
+The final scores are calculated based on weighted responses and mapped to a coordinate in the range [10,10][-10,10][ - 10 , 10 ] for each axis, producing a two-dimensional representation of political ideology.
 Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models.
-The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o [^2] https://openai.com/api, Claude [^3] https://www.anthropic.com, and Gemini 1.5 Pro [^4] https://ai.google.dev/gemini-api/docs.
-The open-source models include: Mistral-7B-Instruct [^5] https://docs.mistral.ai/api/, DeepSeek [^6] https://lambda.ai/, RoBERTa-large [^7] https://huggingface.co/FacebookAI/xlm-roberta-large, RoBERTa-base11, and BERT-large 12.
+The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o, Claude, and Gemini 1.5 Pro.
+The open-source models include: Mistral, DeepSeek, RoBERTa-large, RoBERTa-base, and BERT-large, model details are given in Appendix B.3.
 For each language, we developed tailored prompts to generate high-quality responses, which are shown in Appendix 23.
- [^8] https://huggingface.co/FacebookAI/xlm-roberta-base
-Response Generation and Parameters: To ensure consistency, we fixed generation parameters across all models: $\text{Temperature}=0.5$, Top-$p$ = 1.
-All responses were generated in a zero-shot multilingual setting.
-For stance detection, we used the multilingual mDeBERTa-v3-base model, which supports over 100 languages, including all five under study.
 Stance Scoring and Mapping: To quantify the model’s political stance on each PCT statement, we employ a two-stage computational approach.
-First, we calculate a scalar stance score from classifier confidence values.
+First, we calculate a scalar stance score from classifier confidence values, for more details see Appendix B.4.
 Second, we discretize this continuous score into ordinal categories that align with the PCT response format.
-This allows us to convert raw classifier outputs into interpretable symbolic judgments.
+We discretize stance by selecting the class with the highest softmax confidence score among four candidate labels.
+The continuous stance score is then calculated by multiplying the confidence value with a fixed weight: $\pm$10 for Strongly Agree/Disagree and $\pm$5 for Agree/Disagree.
+This approach reflects both the categorical stance and the model’s certainty.
+This allows us to convert raw classifier outputs into interpretable symbolic judgements.
 Let $A_{s},A,D,D_{s}\in[0,1]$ denote the predicted confidence scores corresponding to the classes Strongly Agree, Agree, Disagree, and Strongly Disagree, respectively.
-1. Stance Score Computation: We define a scoring function $f:[0,1]^{4}\rightarrow[-10,10]$ to assign a value based on the dominant confidence class:
+1. Stance Score Computation: We define a scoring function $f:[0,1]^{4}\rightarrow[-10,10]$ to assign a value based on the dominant class:
 $f(A_{s},A,D,D_{s})=\begin{cases}10\cdot A_{s}&\text{if }A_{s}>\max(A,D,D_{s})\\ 5\cdot A&\text{if }A>\max(A_{s},D,D_{s})\\ -10\cdot D_{s}&\text{if }D_{s}>\max(D,A,A_{s})\\ -5\cdot D&\text{otherwise}\end{cases}$
@@
 Stance Discretization: To facilitate comparative analysis across models and statements, we define a discretization function $g:[-10,10]\rightarrow\{0,1,2,3\}$ that maps the continuous score to categorical labels using a symmetric threshold parameter $\tau>0$:
-$g(S)=\begin{cases}3&\text{if }S\geq 2\tau\\ 2&\text{if }0\leq S<2\tau\\ 1&\text{if }-2\tau<S<0\\ 0&\text{if }S\leq-2\tau\end{cases}$
-This results in an ordinal stance label interpreted as:
-  • 3 – Strongly Agree
-  • 2 – Agree
-  • 1 – Disagree
-  • 0 – Strongly Disagree
-By mapping soft classifier outputs to these well-defined categories, we ensure that downstream aggregation and political leaning visualization remain interpretable and robust.
-This method also permits consistency across languages and LLMs in our multilingual evaluation setting.
-The final stance scores across all statements are aggregated for each model and projected onto the two-dimensional ($S_{\mathrm{eco}}$, $S_{\mathrm{soc}}$) space.
-This facilitates a structured evaluation of political alignment and model behavior across both ideological dimensions.
+$g(S)=\begin{cases}3&\text{if }S\geq 2\tau\\ 2&\text{if }0\leq S<2\tau\\ 1&\text{if }-2\tau<S<0\\ 0&\text{if }S\leq-2\tau\end{cases}$ (1)
+[…]
 3.2 Ideological Framing Analysis
@@
 (2023), Language Policy, Welfare and Charity, Religious Minorities Rights, Policing and Surveillance.
-Each topic was translated into five low-resource Pakistani languages to enable multilingual evaluation.
+Each topic was translated into five Pakistani languages to enable multilingual evaluation.
 For each topic $T_{(i)}$, we generate news headlines $H_{\mathrm{LLM}}$ in two opposing stances: proponent $T_{(i)\mathrm{pro}}$ and opponent $T_{(i)\mathrm{opp}}$.
@@
 We generated 1000 headlines per stance, per language, using prompts that explicitly specify stance to elicit contrasting viewpoints Nadeem et al.
-(2021) (see Appendix  15 for prompt details).
-Frame Dimension Classification: To investigate the ideological narratives, we classify generated headlines using Boydstun’s 15 cross-cutting frame dimensions Boydstun et al.
-(2014), which capture salient political discourse themes such as “Economics,” “Morality,” “Health and Safety,” and “Cultural Identity.” These topic-agnostic frames facilitate comparative analysis across models and topics Hamborg (2020), for details, see Appendix D.
-Frame classification was performed using GPT-3.5-turbo with bilingual prompts, computing frame ratios to identify dominant framing strategies employed by each model Ziems and Yang (2021).
-Entity-Based Framing Analysis: Beyond content frames, we analyze framing at the entity level by extracting and quantifying frequently mentioned named entities (e.g., political figures, countries, organizations) from generated headlines.
+(2021) (see Appendix 23 for prompt details and Appendix B.2for prompting strategy and reproducibility details).
+Frame Dimension Classification: To examine ideological narratives in model-generated content, we classify headlines using Boydstun’s 15 cross-cutting frame dimensions (Boydstun et al., 2014), which encompass salient themes such as Economics, Morality, Health and Safety, and Cultural Identity.
+These topic-independent frames enable consistent comparative analysis across models and topics (Hamborg, 2020).
+We employ GPT-3.5-turbo with bilingual prompts to classify each headline into one or more frames, enhancing contextual understanding in Pakistani languages.
+For each topic–stance pair $(t,s)$, we compute the frame ratio for frame $f_{i}$ as:
+$\text{FrameRatio}_{t,s}(f_{i})=\frac{c_{t,s}(f_{i})}{N_{t,s}}$ (2)
+[…]
+Entity-Based Framing Analysis: To explore how models frame specific actors or institutions, we extract named entities—such as political figures, countries, and organizations—from Urdu headlines using a multilingual NER model.
+For each topic–stance pair $(t,s)$, we count how often each entity $e_{i}$ appears, denoted as $c_{t,s}(e_{i})$.
+We then calculate its relative Prominence P using:
+$\text{P}_{t,s}(e_{i})=\frac{c_{t,s}(e_{i})}{\sum_{j=1}^{n}c_{t,s}(e_{j})}$ (3)
+This score reflects which entities are most emphasized in model outputs, offering insights into how narratives center around particular individuals or groups.
 Frequent entity mentions serve as a subtle framing device, indicating which actors or groups models emphasize Devlin et al.
@@
 NER was conducted using a pretrained bert-base-multilingual-cased model  [^4] https://huggingface.co/google-bert/bert-base-multilingual-cased.
-Sentiment Polarity Towards Entities: To assess stylistic and attitudinal bias, we analyze the sentiment polarity expressed towards entities within headlines.
-Using a fine-tuned XLM-RoBERTa model for Urdu sentiment analysis Fan et al.
-(2019), we categorize sentiment as positive, negative, or neutral Saez-Trumper et al.
+[…]
+$\Pr^{(k)}_{t,s}(e_{i})=\frac{S_{e_{i}}^{(k)}}{\sum_{k^{\prime}}S_{e_{i}}^{(k^{\prime})}}$ (4)
+We define the dominant sentiment polarity for entity $e_{i}$ as:
+$\text{SentimentBias}_{t,s}(e_{i})=\arg\max_{k}\left(\Pr^{(k)}_{t,s}(e_{i})\right)$ (5)
+These scores expose entity-level sentiment bias patterns across topics and stances (Spliethöver et al., 2022; Roy and Goldwasser, 2020).
+4 Experimental Settings
+Dataset: We introduce and release a novel multilingual dataset designed to support political bias analysis across five Pakistani languages: Urdu, Punjabi, Sindhi, Balochi, and Pashto [^5] https://anonymous.4open.science/r/PoliticalBiasEvaluation-10DE.
+The dataset comprises two main components:
+  • Political Compass Test (PCT) Translations: This segment includes 62 culturally adapted political statements translated into each of the five target languages as detailed Table in Appendix 28.
+Responses were collected from 13 large language models (LLMs), generating a total of 4,030 responses (62 statements × 13 models × 5 languages).
+we employed three native speakers per language with triple verification to ensure full linguistic and semantic fidelity.
+Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus.
+  • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see details in Appendix 6.
+Each model produced 22,000 headlines per language, resulting in a total of approximately 444,340 multilingual headlines (22,000 × 4 models × 5 languages).
+A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement and semantic consistency reflecting moderate inter-annotator agreement given the subjective nature of generated content, as Shown in Figure 3.
+The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation.
+All annotators were recruited from linguistically representative Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language.—and were compensated fairly.
+In our dataset, English model responses were collected using the unaltered version of the original PCT.
+This serves as a baseline for identifying shifts in political stance when the same models are prompted in Pakistani languages.
+Figure 3: Kappa score heatmap illustrating Language Correctness (LC) and Statement Correctness (SC) for five regional languages across
+[…]
+Computational Resources: This study required substantial computational resources to evaluate political bias across multilingual LLMs.
+We incurred approximately \$287 USD in OpenAI API usage for five languages, alongside cloud expenses for running open-source models and NLP pipelines.
+The total budget was around \$350 USD.
+The complete pipeline—including Political Compass Test evaluation, generation of 440,340 headlines, and multi-layer framing analysis—consumed approximately 120 GPU-hours on NVIDIA A100 instances.
+5 Results and Analysis
+5.1 Political Stance Distribution Across Languages
+Figure 4 presents political stance outcomes from the Political Compass Test (PCT) across five Pakistani languages.
+Most LLMs cluster in the libertarian-left quadrant, reflecting progressive economic and independent social values.
+Claude shows the strongest libertarian stance, while GPT-4-turbo leans most economically left.
+Models like GPT-3.5-turbo and OpenAI o1-mini shift toward authoritarian-right in Urdu, highlighting language-specific influences.
+Sindhi remains consistently libertarian-left, whereas BERT variants lean right across languages.
+GPT models trend authoritarian-left in regional contexts, unlike open-source models which remain more liberal-libertarian.
+These findings underscore the need for multilingual political bias evaluation to ensure culturally equitable model behavior Johnson and Goldwasser (2016).
+Figure 4: Political leaning of open source and closed source models used for Pakistani language shows diverse inclination across LLM
+Cross Language Bias Patterns: Our findings highlight significant cultural variation in LLM behavior.
+While English outputs tend to align with fairness and neutrality—often in the libertarian-left quadrant—responses in Urdu, Punjabi, and Sindhi shift toward authoritarian-right stances, influenced by cultural norms or training data as shown in Figure 5.
+Pashto remains closest to English in political leaning.
+Even fine-tuned Urdu models retain or amplify these biases (Figure 13).
+Error analysis (Figure 16) shows English as the most stable, whereas Pakistani languages exhibit greater variance and bias.
+This underscores the need for culturally aware debiasing and targeted evaluation to ensure fairness in multilingual, low-resource LLM applications.
+Figure 5: Deterministic decoding reveals political bias shifts in LLMs across six Pakistani languages and English, exposing deep cross-linguistic bias under deterministic conditions.
+Figure 6: Entity-level analysis of Urdu language outputs across LLMs.
+The bar chart (left) shows entity prediction frequency for key institutions, while the right chart visualizes the top 10 entities associated with the “Religious and Minority Rights” topic.
+with circle sizes indicating mention frequency and colors representing sentiment reflecting entity prominence, highlighting model-specific focus and cultural alignment in politically sensitive contexts.
+Cultural Adaptive Models are Less Biased: We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs, to analyse it, we fine-tuned models: Bert-base [^6] https://huggingface.co/google-bert/bert-base-cased,Gemma7b [^7] https://huggingface.co/google/gemma-7b, GPT-3.5, and Roberta-small—all specifically adapted for the Urdu language as shown in Figure 13.
+The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$.
+Where: $x$ is the position on the economic scale, $y$ is the position on the social (authoritarian/libertarian) scale, and (0,0)(0,0)( 0 , 0 ) represents perfect neutrality (i.e., bias score = 0).
+Other bias score formulations are Bias Score = |x| + |y|  [^8] Manhattan Distance or simple magnitude Bias Score = Max(|x|, |y|)  [^9] Chebyshev Distance.
+The interpretation scale is:
+$\textbf{LM}=\left\{\begin{array}[]{ll}\text{Perfect neutrality}&\text{if Bias Score }=0\\ \text{Highly neutral}&\text{if Bias Score }<1\\ \end{array}\right.$
+Model classification by bias score is measured as:
+$\textbf{LM }=\left\{\begin{array}[]{ll}\text{Neutral model}&\text{if Bias Score }\in[0,1)\\ \text{Biased model}&\text{if Bias Score }\geq 1\\ \end{array}\right.$
+Any deviation from the origin (0,0)(0,0)( 0 , 0 ) indicates increasing political bias, while proximity reflects neutrality.
+Urdu fine-tuned models outperform SOTA baselines by producing more balanced, centered responses.
+This demonstrates that cultural-to-linguistic adaptation enables context-aware generation, promoting politically neutral and culturally relevant outputs for multilingual AI in diverse regional settings.
+5.2 Framing Bias Analysis
+Content Bias: In content bias analysis, we evaluated framing dimension and entitiy frequency analysis, which are described below.
+Framing Dimension: All models exhibit varying uses of cultural identity frames when discussing religious minority issues in Pakistan as shown in Figure 12.
+Claude emphasizes universalist fairness, while GPT-4 integrates fairness, morality, and regulation.
+For the death penalty, Claude and Gemini rely on morality, whereas GPT-4 and DeepSeek favor policy frames.
+Shared use of constitutional framing reflects Islamic legal influence.
+Abstract frames lead to higher model error rates, as shown in Figure 7 Fazal (2022).
+Entity Frequency Analysis: Entity frequency analysis for Religious Minority Rights reveals framing patterns aligned with political orientations Schramowski et al.
 (2022).
-Aggregating sentiment across samples yields a profile of the model’s overall stance tone towards each topic, revealing subtle stylistic bias Spliethöver et al.
-(2022); Roy and Goldwasser (2020).
-4 Experimental Settings
-Dataset: We introduce and release a novel multilingual dataset designed to support political bias analysis across five low-resource Pakistani languages: Urdu, Punjabi, Sindhi, Balochi, and Pashto [^12] https://anonymous.4open.science/r/PoliticalBiasEvaluation-10DE.
-The dataset comprises two main components:
-  • Political Compass Test (PCT) Translations: This segment includes 62 culturally adapted political statements translated into each of the five target languages.
-Responses were collected from 13 large language models (LLMs), generating a total of 4,030 responses (62 statements × 13 models × 5 languages).
-All translations were produced and validated by native speakers, with triple verification to ensure full linguistic and semantic fidelity.
-Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus.
-  • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see detail in Appendix 2.
-Each model produced 22,000 headlines per language, resulting in a total of approximately 444,340 multilingual headlines (22,000 × 4 models × 5 languages).
-A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement rate of 98.5%.
-The corresponding Fleiss’ kappa score for semantic consistency in the DeepSeek model was 0.34, reflecting moderate inter-annotator agreement given the subjective nature of generated content.
+Figure 6 shows that models exhibiting more authoritarian stances tend to frame minority rights through institutional and geographic hierarchies, emphasizing state and regional across top 10 entities, such as Pakistan, the Supreme Court, and regional bodies serve to situate the discourse within legal and geographic contexts, words like rights and law reinforce a rights-based framing.
+Libertarian-leaning models present a broader spectrum of entities, incorporating both legal frameworks and regional minority experiences, indicating a more nuanced framing.
+Stylistic Bias: Lexical polarity analysis highlights how LLMs stylistically frame within Pakistani political discourse.
+Figure 14 shows some of society’s most sensitive and debated issues.
+Same-sex marriage rights stand out with the most polarized coverage—yet interestingly, the sentiment leans slightly positive, hinting at a complex and emotionally charged discourse.
+Language policy, welfare and charity, and education follow closely, marked by passionate debate but generally hopeful tone.
+On the other hand, deeply rooted religious and moral issues like abortion, blasphemy laws, and the death penalty show intense division and overwhelmingly negative sentiment.
+When it comes to government performance, coverage tends to be both critical and sharply divided—reflecting growing public frustration as shown in Figure 17.
+Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression.
+For more analysis and results see Appendix C.
+Figure 7: Error rates across political discourse dimensions reveal how confidently models handle different types of content.
+6 Findings
+6.1 Model-Level Bias Interpretation
+Figure  8 highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini.
+The results highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini.
+To further interpret how political bias manifests in model behaviour, we conducted a detailed, multidimensional analysis of the DeepSeek model’s visual and quantitative breakdowns, which are provided in the Appendix Figure 21.
+Figure 8: Lexical Polarity Rate (LPR) comparison across 11 sociopolitical topics for LLMs.
+6.2 Cross-Cultural model insights
+Our analysis reveals distinct model behaviors across cultural-linguistic contexts.
+GPT models show high cultural sensitivity, shifting from libertarian-left in English to authoritarian-left in Pakistani languages (Figure 4).
+Claude remains consistently libertarian with progressive sentiment.
+Open-source models vary—Deep Seek offers stable entity selection, while Gemini shows notable ideological shifts.
+These findings challenge universal bias mitigation, emphasizing that effective global AI requires balancing fairness with cultural nuance—something current models often fail to achieve.
+6.3 Multilingual Influence on Political Bias
+The clustering plots of political stance detection across languages reveal that LLMs exhibit biased behavior influenced by political and cultural context.
+This variation highlights imbalances in training data, with English-dominant models adapting differently in low-resource languages.
+Figure 18 shows consistent support for topics like Education but sharp divergence on Language Policy.
+These findings suggest that multilingualism can generate language-specific political personalities, raising concerns that global LLM deployment may unintentionally reinforce cultural biases depending on the language used.
+7 Conclusion
+This study presents the first large-scale investigation of political bias in large language models across five low-resource yet widely spoken Pakistani languages: Urdu, Punjabi, Sindhi, Pashto, and Balochi.
+We introduce a novel, culturally adapted evaluation framework that integrates ideological stance scoring with discourse-level framing analysis—capturing both what the model says and how it says it.
+Our findings uncover systematic, language-conditioned shifts in political bias, demonstrating how linguistic and cultural context significantly shapes model behavior.
+By bridging the Political Compass Test (PCT) with narrative framing dimensions, our methodology offers a reproducible and extensible approach for diagnosing political bias in multilingual LLMs.
+This work fills a critical gap in multilingual NLP and contributes a practical auditing tool for building culturally grounded, fair, and context-aware AI systems.
+Limitations
+This study has several methodological and scope-related limitations.
+Our analysis is restricted to five Pakistani languages, omitting other regional or global languages that may exhibit different ideological patterns.
+We also exclude high-resource languages like French or Arabic, which could have served as cultural counterfactuals to help distinguish between linguistic and cultural influences in LLM pretraining.
+While we adapt the Political Compass Test (PCT), its original design is rooted in Western political thought and may not fully capture culturally specific structures like Islamic jurisprudence or tribal governance.
+Additionally, our sentiment and entity analysis tools are primarily trained on English, potentially reducing accuracy and missing localized expressions.
+We do not explore prompt-induced framing bias, which may affect stance independently of model ideology.
+Finally, the static nature of training data limits our ability to assess real-time political shifts.
+Future work should explore culturally sensitive bias mitigation for low-resource settings.
+Ethical Statement
+This research was conducted with strict devotion to ethical principles, ensuring cultural sensitivity and participant welfare.
+Content generation carefully avoided potentially harmful or inflammatory material while maintaining analytical integrity.
+We acknowledge possible biases in our Western-developed evaluation frameworks and commit to transparent reporting of limitations.
+The dataset excludes personally identifiable information and extreme political content that could incite violence or discrimination.
+We recognize the responsibility of AI bias research in multicultural contexts and emphasize that our findings should inform inclusive AI development rather than reinforce stereotypes.
+This work aims to promote reasonable AI systems that respect diverse political perspectives and cultural values across Global South communities.
+Appendix A Appendix
+A.1 Dataset Contribution
+Our research provides a valuable multilingual dataset that spans five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies.
+For the Political Compass Test, the statement are translated into five languages shown in the Table 28.
+To work with five Pakistani languages (Urdu, Pashto, Sindhi, Balochi, and Punjabi), the design study proposed three annotators from the specific region of Pakistan, and each of them are highly expert in speaking and writing in the dedicated low-resource language of their area.
 The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation.
-All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly.
+We chose the Method: Triple-verified by native speakers.
+All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions.
+The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  9.
+The generated response on 62 culturally adapted political statements in all five languages as shown in Table 28, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 3.
+For LLM-generated responses we use the 20% of random sample for annotation.
+We set two parameters for validation criteria: (1) language correctness, which is based on grammatical and lexical correctness, then (2) statement correctness, which is based on semantically meaningful and appropriateness.
+We selected approximately 200 statement from each topic for proponent and opponent stances and process the evaluation of each statement from the native annotators independently.
+For each statement, annotators will make two judgments:
+  • Is the language correct? (Yes/No $\rightarrow$ 1/0)
+  • Is the statement grammatically meaningful/sensible? (Yes/No $\rightarrow$ 1/0)
+A.2 Language Translation Procedure
+For each of the five target languages, prompts were translated from English by a team of three bilingual native speakers.
+This was followed by model generation in the respective language, forming a two-step pipeline.
+While multi-translator input reduced individual bias, we acknowledge the absence of back-translation or inter-annotator agreement as a limitation that may introduce subtle framing or cultural interpretation bias in multilingual political contexts.
+A.3 Translation Assignment Procedure
+Each statement was independently translated into the target language by three bilingual native speakers.
+Translations were then compared, and a final version was selected through majority agreement or collaborative consensus when needed.
+This approach ensured semantic accuracy while reducing individual translator bias.
+For the annotation we created a template for rating between 0 and 1, where 0 is for NO and 1 is for YES.
+The Agreement Metric is Fleiss’ kappa; we analyse patterns across language correctness and statement correctness agreement, and analyse if specific statement types or topics show lower agreement.
+The interpretations are based on standard kappa ranges, that is:
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+The Fleiss’ kappa statistic is calculated as:
+$\kappa=\frac{P-P_{e}}{1-P_{e}}$
+where $P$ represents the observed agreement between annotators, $P_{e}$ represents the agreement expected by chance, and $\kappa$ ranges from -1 to 1.
+Figure 9: Heatmap of the pct statement translations on all five languages
+The significance of calculating the kappa score for Pakistani languages are that how efficiently LLM performs in different languages with multiple resources and establishes benchmarks for future work in Pakistani language NLP with ca omprehensive approach to evaluating the reliability of your LLM-generated content across these five Pakistani languages.
+Our research provides a valuable multilingual dataset contribution across five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies.
 Category Details
@@
 Table 1: Summary Statistics for Political Compass Test Translations and Headline Generation Corpus
-Computational Resources: This study demanded significant computational resources to evaluate political bias through multiple language models and linguistic contexts.
-We utilized commercial API services for model access, with OpenAI API costs approximately \$287 USD for open-AI model evaluations (GPT-3.5-turbo, GPT-4, GPT-4-turbo, open-ai mini, open-ai preview) in five languages.
-Additional computational expenses included cloud computing resources for running open-source models and multilingual NLP pipelines for entity identification and sentiment analysis.
-The total computational budget, including API costs, cloud infrastructure, and model inference, was estimated at \$350 USD.
-Processing the complete experimental pipeline, including PCT evaluation, 440,340 headline generation, and three-method framing analysis, required approximately 120 GPU-hours on NVIDIA A100 instances.
-5 Results and Analysis
-5.1 Political Stance Distribution Across Languages
-Figure 2 presents political stance detection results for various LLMs evaluated in Urdu, Punjabi, Balochi, Sindhi, and Pashto using the Political Compass Test (PCT).
-The majority of models cluster within the libertarian-left quadrant, indicating progressive economic views coupled with socially independent positions.
-Among them, Claude exhibits the strongest libertarian tendencies, while GPT-4-turbo scores the most pronounced leftward economic stance.
-Interestingly, some models, such as OpenAI o1-mini and GPT-3.5-turbo, position in the authoritarian-right quadrant when responding in Urdu, indicating traditional or conservative leanings in this language context.
-Sindhi results largely reflect libertarian-left stances across models.
-Overall, the PCT results reveal considerable variance in political positioning across models and languages.
-GPT-based models consistently shift toward authoritarian-left stances in regional languages (Urdu, Punjabi, Pashto, Balochi), whereas open-source models like Claude, Gemini, Mistral, and DeepSeek predominantly maintain libertarian-left orientations Johnson and Goldwasser (2016).
-This suggests that GPT models may interpret non-Western linguistic inputs as favoring stronger state control.
-BERT variants generally display right-leaning tendencies irrespective of language.
-These findings underscore how language choice significantly influences LLMs’ expressed political biases, reinforcing the necessity for multilingual evaluations to fully capture bias dimensions.
-Figure 2: Political leaning of open source and closed source models used for low-resource language shows diverse inclination across LLM
-5.1.1 Cultural Adaptive Models are Less Biased
-We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs, to analyse it, we fine-tuned models: Bert-base [^13] https://huggingface.co/google-bert/bert-base-cased,Gemma7b [^14] https://huggingface.co/google/gemma-7b, GPT-3.5, and Roberta-small—all specifically adapted for the Urdu language as shown in Figure 3.
-The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as:
-Bias $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$
-Where:
-  • $x$ = position on the economic scale
-  • $y$ = position on the social (authoritarian/libertarian) scale
-  • $(0,0)$ = perfect neutrality (bias score = 0)
-Other bias score formulations are Bias Score = |x| + |y|  [^15] Manhattan Distance or simple magnitude Bias Score = Max(|x|, |y|)  [^16] Chebyshev Distance. The interpretation scale is:
-Figure 3: Political leaning of four open-source LLMs used for the Fine-tuning on Urdu language: a deep analysis on cultural nuance and political inclination of LLMs
-$\textbf{LM}=\left\{\begin{array}[]{ll}\text{Perfect neutrality}&\text{if Bias Score }=0\\ \text{Highly neutral}&\text{if Bias Score }<1\\ \end{array}\right.$
-Model classification by bias score is measured as:
-LM =$\textbf{LM }=\left\{\begin{array}[]{ll}\text{Neutral model}&\text{if Bias Score }\in[0,1)\\ \text{Biased model}&\text{if Bias Score }\geq 1\\ \end{array}\right.$
-The closer the bias score is to zero, the more politically neutral the model is.
-Any deviation from the origin $(0,0)$ indicates increasing political bias.
-This mathematical framework clearly shows that zero represents perfect neutrality, and any movement away from the center introduces measurable bias.
-The Urdu fine-tuned models demonstrated superior performance in realizing more composed and centered political positioning compared to state-of-the-art (SOTA) models.
-These models effectively reduced political bias and achieved improved neutrality.
-This outcome indicates that targeted fine-tuning on Urdu linguistic features helps models to better engage with nuanced cultural and political discourse.
-These models show knowledgeable and context-aware behavior in terms of political and cultural perceptions rooted specifically in Pakistan.
-The major implication of Fine-tuning for cultural $\rightarrow$ linguistic adaptation can integrate regional political viewpoints, producing models that are culturally aware while remaining politically neutral.
-This highlights the necessary balance between cultural relevance and political neutrality in AI systems.
-Figure 4: Frame dimension ratio for "Religious minority rights" topic for four models. Overall similar but most variance observed in "Morality", "Fairness and Equality" frames.
-Figure 5: Entity-Based Frame Analysis on Pakistan’s Religious Rights.
-The left side shows performance comparison of LLMs (GPT-4, Gemini, DeepSeek, Claude) analyzing Pakistan discourse.
-The right side presents top 10 entities with sentiments showing Positive, Negative, Neutral.
-Supreme Court and National Commission as central nodes.
-Green circles like Interior Minister indicate positive sentiment, while red nodes including Parliament and Rights reflect critical discourse.
-Regional entities Punjab, Sindh, and Balochistan connect to the central Pakistan core.
-Dashed circles represent unique entities.
-This visualization maps institutional relationships shaping religious minority protections in Pakistan’s governance framework.
-5.2 Framing Bias Analysis
-5.2.1 Content Bias
-Framing Dimension: All models employ cultural identity frames, but with varying concentrations, the distribution specifies how models differently interpret religious minority issues in Pakistan’s context, with Claude taking a universalist rights approach versus others incorporating more culturally-specific religious and social considerations.
-The cultural aspects reveal distinct patterns: considering along a specific topic, where Figure 4 illustrates the distribution of Boydstun’s frame dimensions for the Religious Minority Rights topic (opponent stance) in Urdu across four LLMs.
-Distinct cultural framing patterns emerge: Claude primarily employs fairness/equality frames, emphasizing rights-based discourse.
-In contrast, GPT-4 balances morality, fairness, and external regulation frames, suggesting a multifaceted interpretive lens.
-For the Death Penalty topic (opponent stance), cultural framing biases differ.
-Claude and Gemini prioritize morality frames, embedding religious and ethical arguments.
-GPT-4 and DeepSeek focus on policy evaluation and fairness frames.
-Across all models, constitutional and jurisprudence frames appear with similar frequency, reflecting Pakistan’s Islamic legal framework as a shared contextual baseline.
-Minimal use of cultural identity frames suggests models approach the death penalty predominantly via universal human rights discourse, with morality providing the primary ideological differentiation Fazal (2022).
-Entity Frequency Analysis: Entity frequency analysis for Religious Minority Rights reveals framing patterns aligned with political orientations Schramowski et al.
+Appendix B Technical details
+B.1 Stance Detection vs. Our Ideological Mapping
+Standard stance detection typically involves classifying whether a text expresses a pro, con, or neutral position toward a specific target or claim Mohammad et al.
 (2016).
-Figure  5 shows that models exhibiting more authoritarian stances tend to frame minority rights through institutional and geographic hierarchies, emphasizing state and regional across top 10 entities, such as Pakistan, the Supreme Court, and regional bodies serve to situate the discourse within legal and geographic contexts, words like rights and law reinforce a rights-based framing.
-Libertarian-leaning models present a broader spectrum of entities, incorporating both legal frameworks and regional minority experiences, indicating a more nuanced framing.
-Figure 6: Deep Seek model topic by polarization in Urdu Headlines: Topics ranked by opinion extremity, with color dots indicating positive (green) or negative (red) sentiment direction.
-5.2.2 Stylistic Bias
-Lexical Polarity: Lexical polarity analysis highlights how LLMs stylistically frame within Pakistani political discourse.
-Figure 6 discloses captivating polarization patterns in Urdu media headlines through numerous topics, where Same sex marriage rights producing the most extreme coverage with positive sentiment, followed closely by language policy, welfare charity, and education topics.
-Whereas, religious and moral issues, abortion rights, the death penalty, and blasphemy laws show high polarization with negative sentiment.
-Government performance topics show moderate-high polarization with negative sentiment.
-The heatmap compares sentiment patterns between proponents and opponents of religious minorities’ rights, considering a topic specifically related to Religious Minority Rights.
-Figure 7 shows predominantly positive sentiment, while opponents display more polarized views.
-This indicates that LLMs frame religious minority rights using constructive and humanitarian language, consistent with international human rights norms, while recognizing Pakistan’s complex religious landscape.
-Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression.
-Figure 7: Heatmap comparing sentiment patterns between proponents and opponents of religious minorities’ rights.
-Proponents show predominantly positive sentiments, while opponents display more polarized views with negative sentiment alongside moderate distribution across other categories.
-Both groups maintain similar neutral sentiment levels.
-5.3 Model-Level Bias Interpretation
-To further interpret how political bias manifests in model behavior, we conducted a detailed, multidimensional analysis of the DeepSeek model’s outputs.
-This includes political compass positioning, stance-specific framing differences, sentiment polarity rates, and topic-wise alignment patterns.
-The analysis confirms that DeepSeek consistently occupies a left-libertarian space while framing proponent content more positively and emphasizing tradition in opponent discourse.
-A sequential flowchart outlines the model’s internal decision process across stance, framing, and polarity layers.
-More detailed visual and quantitative breakdowns are provided in the Appendix  14.
-5.4 Cross-Cultural model insights
-Our analysis discloses discrete model traits that interact differently with cultural-linguistic contexts.
-Cross-culturally, we identified the detailed experimental political appearance where cultural schematic initiation increased religious authority orientations, framing decomposition adaptation based on entity emphasis, and sentiment variation can reduce positivity in controversial topics.
-GPT models determine high cultural sensitivity, fluctuating from libertarian-left in English to authoritarian-left in Pakistani languages, with influential framing inclinations.
-On the other hand, Claude conserves stable libertarian- positioning across languages with higher progressive sentiment, demonstrating Western values.
-Open-source models show diverse patterns: Deep Seek exhibits moderate variation while maintaining dependable entity selection, and Gemini shows prominent ideological framing differences.
-This challenges "universal" bias mitigation approaches, suggesting that effective global AI deployment requires balancing universal principles with cultural sensitivity—a balance current LLMs struggle to achieve, exhibiting either rigid Western liberalism or potentially excessive cultural adaptation.
-5.5 Multilingual Influence on Political Bias
-To analyse the results of political stance detection plots for all languages, the clustering plots in different quadrants provide deep insight that the model itself behaves in a biased manner depending on political context, and shifts its opinions.
-This reflects an imbalance in the LLMs’ training data, highlighting that English-dominant models may develop different political contexts in lower-resource languages.
-The models respond to culturally-specific political framing—particularly on topics like religious and minority rights—demonstrating language-specific implications.
-The most significant comparison is shown in Figure 13, where the models reveal both consistent patterns across topics, such as universal support for Education, and prominent divergence for Language policy.
-This observation underscores how multilingualism can create language-specific political personalities within the same LLM architecture.
-Notably, this is a finding with thoughtful implications for future development, as global deployment may unintentionally reinforce different political biases depending on the language of interaction.
-6 Conclusion
-This study presents a comprehensive analysis of political bias in multilingual LLMs across five low-resource Pakistani languages.
-Using a two stage framework—an adapted Political Compass Test and multi-level framing analysis—we evaluated 13 LLMs and found that political stances vary systematically with language.
-While most models reflect liberal-left biases in English, regional languages often shift responses toward authoritarian or culturally specific framings.
-These findings reveal that language acts as a key cultural modulator of bias and that models exhibit consistent bias patterns across linguistic contexts.
-Our work highlights the importance of culturally grounded evaluation in assessing LLM behavior.
-Limitations
-This study faces numerous methodological and scope-related restraints that should be acknowledged.
-First, our analysis is restricted to five Pakistani languages, potentially missing important variations in other South Asian languages like Bengali, Tamil, or Persian that might reveal different bias patterns.
-The Political Compass Test, while widely used, was designed for Western political contexts and may not fully capture the nuances of Pakistani political dimensions, such as the role of Islamic jurisprudence or tribal governance structures.
-Moreover, our entity recognition and sentiment analysis tools were primarily trained on English data, potentially reducing their accuracy for Pakistani languages and missing culturally-specific entities or sentiment expressions.
-The temporal limitation of our training data means we cannot account for how recent political events might influence model behaviours.
-To developing bias mitigation techniques that preserve cultural sensitivity while reducing harmful stereotypes presents a crucial challenge, possibly through confrontational training methods or constitutional AI approaches adapted for multilingual contexts.
-Ethical Statement
-This research was conducted with strict devotion to ethical principles, ensuring cultural sensitivity and participant welfare.
-Content generation carefully avoided potentially harmful or inflammatory material while maintaining analytical integrity.
-We acknowledge possible biases in our Western-developed evaluation frameworks and commit to transparent reporting of limitations.
-The dataset excludes personally identifiable information and extreme political content that could incite violence or discrimination.
-We recognize the responsibility of AI bias research in multicultural contexts and emphasize that our findings should inform inclusive AI development rather than reinforce stereotypes.
-This work aims to promote reasonable AI systems that respect diverse political perspectives and cultural values across Global South communities.
-Appendix A Appendix
-Dataset Contribution Our research provides a valuable multilingual dataset that spans five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies.
-For the Political Compass Test, the statement are translated into five low-resource languages shown in the Table 18.
-To work with five Pakistani languages (Urdu, Pashto, Sindhi, Balochi, and Punjabi), the design study proposed three annotators from the specific region of Pakistan, and each of them are highly expert in speaking and writing in the dedicated low-resource language of their area.
-The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation.
-We chose the Method: Triple-verified by native speakers.
-All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions.
-The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  8.
-The generated response on 62 culturally-adapted political statements in all five languages as shown in Table 17, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 9, 10, 11, 12.
-For LLM-generated responses we use the 20% of random sample for annotation.
-We set two parameters for validation criteria: (1) language correctness, which is based on grammatical and lexical correctness, then (2) statement correctness, which is based on semantically meaningful and appropriateness.
-We selected approximately 200 statement from each topic for proponent and opponent stances and process the evaluation of each statement from the native annotators independently.
-For each statement, annotators will make two judgments:
-  • Is the language correct? (Yes/No $\rightarrow$ 1/0)
-  • Is the statement grammatically meaningful/sensible? (Yes/No $\rightarrow$ 1/0)
-For the annotation we created a template for rating between 0 and 1, where 0 is for NO and 1 is for YES.
-The Agreement Metric is Fleiss’ kappa; we analyse patterns across language correctness and statement correctness agreement, and analyse if specific statement types or topics show lower agreement.
-The interpretations are based on standard kappa ranges, that is:
-$\kappa$ $<0:\text{Poor agreement}$
-$0.01\leq\kappa$ $\leq 0.20:\text{Slight agreement}$
-$0.21\leq\kappa$ $\leq 0.40:\text{Fair agreement}$
-$0.41\leq\kappa$ $\leq 0.60:\text{Moderate agreement}$
-$0.61\leq\kappa$ $\leq 0.80:\text{Substantial agreement}$
-$0.81\leq\kappa$ $\leq 1.00:\text{Almost perfect agreement}$
-The Fleiss’ kappa statistic is calculated as:
-$\kappa=\frac{P-P_{e}}{1-P_{e}}$
-where $P$ represents the observed agreement between annotators, $P_{e}$ represents the agreement expected by chance, and $\kappa$ ranges from -1 to 1.
-The significance of calculating the kappa score for low-resource languages are that how efficiently LLM performs in different languages with multiple resources and establishes benchmarks for future work in Pakistani language NLP with ca omprehensive approach to evaluating the reliability of your LLM-generated content across these five Pakistani languages.
-Our research provides a valuable multilingual dataset contribution across five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies.
-Figure 8: Heatmap of the pct statement translations on all five languages
-Figure 9: Heatmap of the news headlines generated by Deepseek model across 20% samples. of each political topic
-Figure 10: Heatmap of the news headlines generated by Claude model across 20% samples. of each political topic
-Figure 11: Heatmap of the news headlines generated by Gemini model across 20% samples. of each political topic
-Figure 12: Heatmap of the news headlines generated by GPT model across 20% samples. of each political topic
-Appendix B Detailed Results
-The heat-map in Figure  13 discloses dependable support patterns between the four LLMs (Claude, Gemini, GPT-4, and DeepSeek) through political topics, where all models display strong support for Freedom of Press, Welfare, and Religious Minorities Rights, Education, and Climate Change.
+These approaches focus on binary or ternary stance concerning an explicit target, often in single-turn texts such as tweets.
+In contrast, our methodology uses an adapted Political Compass Test to infer a model’s position in a two-dimensional ideological space (economic and social axes).
+Rather than target-specific classification, we aggregate stance scores across 62 political statements to construct a holistic ideological profile per model and language.
+This offers a structured lens into political bias beyond isolated stance decisions.
+Moreover, we complement this scalar stance mapping with rhetorical framing analysis to examine how ideological leanings are expressed stylistically and narratively—going beyond traditional stance detection’s limited focus on polarity or agreement.
+Stance detection is a widely studied task in NLP Gorrell et al.
+(2019), typically formulated as predicting whether a speaker is in favor or against a known target or topic.
+These tasks are often applied to tweets, debates, or news articles.
+In contrast, our use of the Political Compass Test allows for continuous stance scoring across a spectrum of ideologically salient statements, enabling two-dimensional mapping of model behavior.
+This richer representation is particularly important for analyzing latent political bias in generative LLMs, where stance is not tied to a single topic but emerges across diverse ideological domains.
+B.2 Prompting Strategy and Reproducibility
+We adopt a zero-shot prompting strategy for all LLM generations to avoid bias from exemplars.
+For each of the 62 PCT statements, models are asked to respond in the specified language using a Likert-style agreement format.
+For framing analysis, we generate 1,000 headlines per stance–topic pair using bilingual prompts that explicitly request either supportive or opposing tone (see Appendix 21).
+Frame classification is conducted using GPT-3.5-turbo with a fixed schema based on Boydstun’s taxonomy.
+Named entity and sentiment analyses are automated via mDeBERTa and XLM-RoBERTa.
+All code and prompts are shared publicly for reproducibility.
+B.3 Model Details
+Table 2 provides an overview of the language models used in our bias evaluation.
+It includes both closed-source and open-source models, detailing their type, estimated parameter sizes, and architectures.
+Hyperlinked model names direct to official documentation or repositories, enabling transparency and reproducibility for further comparative analysis.
+Model Architectures and Tuning Details Most of the models we evaluate, including GPT-4, Claude, Gemini, and DeepSeek, are not raw pretrained models but represent fully developed systems with instruction tuning and safety alignment, often including RLHF.
+Our focus is on assessing bias as it appears in real-world, user-facing outputs.
+We include both decoder-based models (e.g., DeepSeek-Chat) and encoder-based models (e.g., BERT, RoBERTa) to explore architectural effects.
+While decoder models are prompted generatively, encoder models are probed using classification on the same inputs.
+We acknowledge that this mix, along with post-training layers, may influence results and that further work is needed to isolate these factors.
+Table 2: Overview of Language Models Used in Bias Evaluation
+Model Name Type Parameters Architecture
+GPT-3.5-turbo Closed-source $\sim$175B (est.) Decoder
+GPT-4-turbo Closed-source $\sim$1.8T (est.) Decoder
+GPT-4 Closed-source $\sim$1.8T (est.) Decoder
+GPT-4o Closed-source $\sim$1.8T (est.) Decoder
+OpenAI o1-mini Closed-source Unknown Decoder
+OpenAI o1-preview Closed-source Unknown Decoder
+Claude-3-Haiku-202403 Closed-source $\sim$13B (est.) Decoder
+Gemini-1.5-Pro Closed-source Unknown Decoder
+Mistral-7B-Instruct-v0.2 Open-source 7B Decoder
+DeepSeek-Chat Open-source 7B Decoder
+XLM-RoBERTa-large Open-source 550M Encoder
+BERT-large Open-source 340M Encoder
+XLM-RoBERTa-base Open-source 270M Encoder
+B.4 On Bias and Limitations of mDeBERTa
+While mDeBERTa-v3-base-mnli-xnli offers strong cross-lingual performance for zero-shot stance classification, we acknowledge that it may carry latent biases inherited from its training on the XNLI corpus.
+The XNLI dataset is primarily derived from translations of English data and may over-represent high-resource languages and Western discourse norms.
+This could influence how disagreement or ambiguity is expressed in lower-resource languages like Urdu or Balochi.
+Although mDeBERTa outperformed alternatives such as XLM-R and mBERT in pilot tests for our target languages, we note that future work should explore culturally fine-tuned models or adversarial probing to better surface language-specific classification bias.
+Role of mDeBERTa in Stance Classification.
+To perform stance classification over multilingual PCT responses, we utilized mDeBERTa-v3-base, a multilingual variant of DeBERTa pretrained on XLM-R corpora, which offers enhanced cross-lingual representation capabilities.
+We selected mDeBERTa over alternatives such as XLM-RoBERTa and mBERT due to its superior performance in zero-shot stance and sentiment classification benchmarks, especially for underrepresented languages.
+Its disentangled attention mechanism and language-agnostic pretraining make it a suitable choice for capturing ideological nuance across the five Pakistani languages evaluated.
+We fine-tuned mDeBERTa on translated PCT examples and constrained the output to four stance labels (Strongly Agree, Agree, Disagree, Strongly Disagree).
+While mDeBERTa performs competitively, we acknowledge potential limitations from English-centric pretraining that may introduce biases or misalignments in culturally specific contexts, which we mitigate through triple-verified human translations and language-specific prompt tuning.
+Future work can explore culturally grounded multilingual encoders tailored to South Asian political discourse.
+B.5 Stance Classification Details
+We use the mDeBERTa-v3-base-mnli-xnli model from HuggingFace’s Transformers library as a zero-shot classifier to assign stance labels.
+The model is prompted with concatenated input: the PCT statement and model response.
+It returns softmax confidence scores across four labels: Strongly Agree, Agree, Disagree, and Strongly Disagree.
+The label with the highest confidence is selected as the predicted stance.
+We compute a numeric stance score by multiplying the winning label’s score by ±5 or ±10.
+For example, a confidence of 0.86 on Strongly Disagree yields a stance score of –8.6.
+This scoring method provides interpretable agreement strength in both ordinal and continuous forms.mDeBERTa-v3-base-mnli-xnli was chosen for stance classification because of its zero-shot multilingual capabilities and language coverage, outperforming alternatives like XLM-R or monolingual classifiers in cross-lingual consistency.
+B.6 Standard PCT Automation
+We follow the original two-dimensional structure of the Political Compass Test, consisting of the economic (Seco) and social (Ssoc) axes.
+Rather than replicate its internal scoring algorithm, we simulate user input by mapping model-generated stance scores to the 4-option Likert scale and input them programmatically into the official PCT interface using Selenium.
+This yields authentic quadrant-level coordinates directly from the source.
+B.7 Language Focus and Fine-tuning Scope.
+In this study, we did not conduct full model fine-tuning due to resource constraints and instead relied on zero-shot and instruction-tuned responses from existing large language models (LLMs) across multiple languages.
+For controlled probing in low-resource settings, we strategically focused on Urdu to examine how political and economic bias manifests in culturally grounded contexts.
+Urdu was selected based on its linguistic richness, wider resource availability, and its status as the mother tongue of Pakistan.
+B.8 Framing Setup
+We acknowledge that using a single model response per prompt may introduce sampling variance in frame predictions.
+Incorporating majority voting across multiple generations, or ensembling across different models, could improve the robustness of frame assignment and reduce random variability.
+We leave this to future work due to API cost constraints.
+B.9 Sentiment Classifier
+We employ XLM-RoBERTa (XLM-R) for downstream tasks such as multilingual frame classification due to its robust performance across 100 languages, including low-resource ones like Urdu, Punjabi, and Pashto.
+Trained on CommonCrawl data in a self-supervised manner, XLM-R provides strong cross-lingual generalization, making it well-suited for tasks where labeled data is scarce or unavailable in the target language.
+Compared to alternatives like mBERT, XLM-R achieves superior results in cross-lingual transfer, particularly for sentence-level classification tasks, while maintaining consistency across diverse scripts.
+Its architecture also allows effective integration with frame-tagging pipelines in our zero-shot or few-shot evaluations.
+Appendix C Detailed Results
+C.1 Ideological Consistency of PCT Responses
+Figure 10 illustrates the distribution of model responses to Political Compass Test (PCT) prompts projected along the economic (Seco) and social (Ssoc) axes, with bootstrapped 95% confidence intervals shown as translucent error bars.
+Each dot corresponds to a single prompt, and its position reflects the average ideological stance expressed by the model.
+The plot reveals a strong diagonal clustering from the lower-left to the upper-right quadrant, indicating a high correlation between the model’s economic and social leanings.
+Despite variation introduced through multilingual translations, the narrow spread of error bars for most points suggests stable and consistent model behavior.
+A few prompts with wider intervals reflect ideologically ambiguous or culturally sensitive content.
+This visualization offers a fine-grained and interpretable view of model ideology, moving beyond discrete stance labels and enabling deeper insight into alignment patterns across ideological dimensions.
+C.2 Ideological Leanings of LLMs Across Political Topics
+The heat-map in Figure  18 discloses dependable support patterns between the four LLMs (Claude, Gemini, GPT-4, and DeepSeek) through political topics, where all models display strong support for Freedom of Press, Welfare, and Religious Minorities Rights, Education, and Climate Change.
 Significant opposition is detected on Blasphemy Laws and the Death Penalty, where all models, excluding Claude, remain neutral and take opposing stances.
@@
 Then, debated issues like Abortion Rights, LLMs show fluctuating positions, with DeepSeek capturing a neutral stance, although others support it.
+Figure 10: Mean stance positions on economic (Seco) and social (Ssoc) axes for PCT prompts, with 95% bootstrapped confidence intervals ($n=10$).
+The diagonal pattern indicates a strong correlation between ideological dimensions.
+Figure 11: Stance comparison Religious Minorities Rights
+Figure 12: Frame dimension ratio for "Religious minority rights" topic for four models. Overall similar but most variance observed in "Morality", "Fairness and Equality" frames.
+Figure 13: Political leaning of four open-source LLMs used for the Fine-tuning on Urdu language: a deep analysis on cultural nuance of political inclination of LLMs
+Figure 14: Deep Seek model topic by polarization in Urdu Headlines: Topics ranked by opinion extremity, with color dots indicating positive (green) or negative (red) sentiment direction.
+Figure 15: Scatter plot showing how model confidence relates to prediction uncertainty.
+When models are highly confident, their bootstrap error can remain significant.
+The trend line ($y=1.15x-0.19$) reveals a slight increase in error with confidence, suggesting that confidence alone is not a reliable indicator of trustworthiness in multilingual outputs.
+C.3 Boydstun Framing Patterns in Pro vs Opp Across LLMs
+Figure 22 presents a comparative analysis of framing dimensions across key sociopolitical topics using four large language models: Claude, DeepSeek, Gemini, and GPT-4.
+Each subplot corresponds to a specific topic (e.g., LGBTQ rights, climate change, education) and illustrates the distribution of Boydstun framing dimensions (e.g., Economic, Morality, Fairness and Equality, Security and Defence).
+The solid bars represent the percentage of responses invoking each frame, with distinctions made between proponent and opponent stances.
+Clear patterns emerge: “Morality” dominates discussions on blasphemy laws, “Fairness and Equality” is prominent in LGBTQ-related topics, and “Capacity and Resources” frequently appears in education debates.
+These results underscore how framing choices vary not only by model, but also by issue and stance—revealing nuanced ideological tendencies embedded in LLM outputs.
+C.4 Boydstun Framing Model Confidence
+Figure 15 illustrates the relationship between model confidence and the intensity of Boydstun Framing in responses across political discourse.
+Each point represents a model output, plotted by its confidence score (X-axis) and normalized framing intensity (Y-axis), with color indicating the dominance of a particular frame dimension (Frame Percentage).
+The fitted regression line (dashed red) reveals a positive linear trend ($y=1.15x-0.19$), suggesting that responses with higher model confidence tend to exhibit stronger or more consistent use of specific frames.
+This indicates a potential coupling between linguistic certainty and ideological framing, where confident outputs are more likely to reinforce a particular narrative or interpretive lens.
+The trend supports the hypothesis that framing is not incidental, but may become more pronounced when the model generates responses it deems more certain—highlighting the need for deeper scrutiny in high-confidence predictions when auditing bias in LLM outputs.
+C.5 Lexical Polarity
+The Figure 20 compares sentiment patterns between proponents and opponents of religious minorities’ rights, considering a topic specifically related to Religious Minority Rights, as shown in Figure 17.
+It also shows predominantly positive sentiment, while opponents display more polarized views.
+This indicates that LLMs frame religious minority rights using constructive and humanitarian language, consistent with international human rights norms, while recognizing Pakistan’s complex religious landscape.
+This finding highlights a critical limitation in relying solely on confidence scores as indicators of reliability, especially in multilingual settings.
+The presence of significant bootstrap variance even at high confidence levels underscores the need for more robust uncertainty-aware evaluation frameworks.
+C.6 Polarization
+Figure 19 shows that claude demonstrates the most positive lexical framing across topics, while Gemini shows the least.
+In terms of polarization, Gemini exhibits the highest variability in sentiment across topics, whereas GPT-4 maintains the consistent and balanced tone, with the lowest average polarization score.
+The results reveal distinct sentiment and polarization profiles across four language models.
+Claude exhibits the highest overall average polarity (+0.080), followed by GPT-4 (+0.070) and DeepSeek (+0.060), indicating generally positive framing, while Gemini produces the least positive responses (+0.030).
+In contrast, Gemini shows the highest average polarization (0.630), suggesting greater variability and potentially more divisive language across topics.
+DeepSeek (0.610) and Claude (0.550) also display moderate polarization, whereas GPT-4 maintains the lowest polarization (0.510), reflecting more balanced sentiment.
+These results highlight GPT-4 as the most tonally consistent model, while Claude is the most positive and Gemini the most polarizing.
+These results highlight GPT-4 as the most tonally consistent model, while Claude is the most positive and Gemini the most polarizing.
+C.7 Political Bias Insights Through Our Framework
 The Figure 21, a combined bias results analysis figure presents various aspects.
@@
 The middle section explains DeepSeek’s model results for handling political content, differentiating between proponent outputs through frames like innovation and opponent outputs highlighting tradition and morality.
+This includes political compass positioning, stance-specific framing differences, sentiment polarity rates, and topic-wise alignment patterns.
+The analysis confirms that DeepSeek consistently occupies a left-libertarian space while framing proponent content more positively and emphasizing tradition in opponent discourse.
+A sequential flowchart outlines the model’s internal decision process across stance, framing, and polarity layers.
 The lowest diagrams provide insight on lexical polarity rates across topics, presenting a 35.00% positive rate for proponent stances versus -26.00% for opponent positions.
-Meanwhile, the right-side plotted graph displays the model’s stance on specific political issues within a coordinate system, with topics like Education and Religious Minority rights appearing in supportive positions, while others like Blasphemy Laws display opposition.
-Finally, the flow chart clarifies how political stance detection leads to bias measurement through linguistic investigation, illustrating how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.
+Meanwhile, the right-side plotted graph displays the model’s stance on specific political issues within a coordinate system, with topics like education and religious minorities appearing in supportive positions, while others like blasphemy laws display opposition, as shown in Figure 11.
+Finally, the flowchart illustrates how political stance detection leads to bias measurement through linguistic investigation, showing how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.
+C.8 Do Models from the Same Family Exhibit the Same Bias?
+While models within the same architectural family often share foundational characteristics and pretraining objectives, our results reveal that political and framing biases are not strictly consistent across family lines—particularly in multilingual settings.
+For instance, OpenAI’s GPT series (GPT-3.5, GPT-4, GPT-4o) generally aligns with libertarian-left positions in English but exhibits divergent quadrant shifts in Pakistani languages, such as GPT-3.5 adopting a more authoritarian-right stance in Urdu.
+Similarly, although Claude models consistently favor fairness-based frames and exhibit ideological stability across languages, Gemini models show pronounced shifts toward legalistic or conservative frames in religious and social topics.
+These findings suggest that language context, fine-tuning procedures, and task framing significantly mediate the expression of bias, even within the same family.
+Consequently, model family lineage alone cannot reliably predict ideological behavior—highlighting the need for language-specific and context-aware evaluations of LLM fairness.
+C.9 Does Model Size Correlate with Political Neutrality?
+Our analysis suggests that while larger language models (e.g., GPT-4, Claude) tend to produce more consistent and coherent responses across multilingual prompts, they are not inherently more politically neutral.
+For instance, GPT-4 consistently leans libertarian-left in English and retains relatively stable ideological positions across languages compared to its smaller counterpart GPT-3.5, which exhibits greater quadrant drift in regional languages.
+This indicates improved alignment and consistency with scale.
+However, neutrality is not guaranteed—larger models may still amplify sociopolitical patterns embedded in their training data.
+Prior studies Liu et al.
+(2021); Hartmann et al.
+(2023) also observe that larger models often internalize liberal-leaning priors due to dominant trends in web-scale corpora.
+Consequently, model size improves reliability but does not equate to fairness or ideological balance, especially in cross-cultural contexts.
+C.10 Framing and Political Bias in NLP.
+Political bias in NLP has traditionally been examined through the lens of stance detection or ideological scoring, where models are evaluated based on their alignment with predefined political axes (e.g., liberal vs conservative, authoritarian vs libertarian).
+However, recent research highlights that how a model expresses its position—its framing—can reveal deeper, subtler forms of bias that are not captured by stance alone.
+Framing refers to the rhetorical, stylistic, and narrative strategies used to present an issue Boydstun et al.
+(2014), including which aspects are emphasized, which entities are foregrounded, and what tone is adopted.
+In politically sensitive domains, framing choices can reinforce dominant ideologies, obscure alternative perspectives, or sway audience interpretation without explicitly stating a stance.
+Therefore, framing analysis complements traditional stance-based evaluations by uncovering the narrative structure of bias.
+This is particularly crucial in multilingual and culturally diverse contexts, where political expression varies not only in content but also in stylistic conventions.
+Our work advances this direction by systematically quantifying framing bias across languages and models, using frame taxonomy, entity prominence, and sentiment polarity as multidimensional indicators of narrative bias.
+C.11 Liberal-Leaning Tendencies in Political Topics.
+Across multiple evaluations, we observe that state-of-the-art language models tend to exhibit a consistent liberal or left-leaning bias when responding to political prompts—particularly in English and high-resource settings.
+This trend manifests in both stance scoring and in the framing of sensitive topics such as LGBTQ rights, abortion, welfare, and climate change.
+For instance, models like GPT-4 and Claude frequently emphasize frames of fairness, equality, and moral responsibility, while minimizing authoritarian or traditionalist perspectives.
+Such patterns align with prior studies Hartmann et al.
+(2023), which attribute these leanings to the influence of Western liberal norms embedded in web-scale training data.
+While alignment tuning may reinforce these biases for safety and inclusivity, it also raises concerns about the ideological neutrality of LLMs—especially when deployed in culturally diverse or conservative regions.
+Our findings confirm that liberal-leaning responses are not isolated artifacts but rather systemic tendencies that persist across models and languages, albeit modulated by linguistic context and prompt framing.
+Table 3: Bootstrap-based Bias Evaluation Metrics for Political Compass Responses Across Models and Languages.
+Model Language Avg Conf Min Conf Max Conf Low Conf Pred Bootstrap Error 95% CI Conf-weighted Error Weighted Mean
+GPT-4-Turbo English 0.602 0.276 0.960 26/62 (6.5%) $\pm$0.518 [1.304, 2.339] $\pm$2.518 2.567
+Urdu 0.658 0.317 0.978 12/62 (19.4%) $\pm$0.786 [-0.480, 1.092] $\pm$3.630 0.188
+Pashto 0.575 0.297 0.977 23/62 (37.1%) $\pm$0.735 [-0.704, 0.765] $\pm$3.363 -0.042
+Punjabi 0.584 0.284 0.972 21/62 (33.9%) $\pm$0.588 [0.896, 2.072] $\pm$3.066 1.537
+Balochi 0.579 0.353 0.889 16/62 (25.8%) $\pm$0.668 [0.176, 1.512] $\pm$2.954 0.877
+Sindhi 0.579 0.314 0.931 19/62 (30.6%) $\pm$0.670 [-0.266, 1.075] $\pm$3.191 0.390
+Gemini-1.5-Pro English 0.836 0.577 0.975 0/62 (0.0%) $\pm$0.898 [-2.394, -0.598] $\pm$4.015 -1.331
+Urdu 0.578 0.317 0.978 25/62 (40.3%) $\pm$0.649 [-2.144, -0.845] $\pm$2.803 -1.748
+Pashto 0.560 0.290 0.985 29/62 (46.8%) $\pm$0.607 [-1.679, -0.466] $\pm$2.977 -1.407
+Punjabi 0.520 0.275 0.916 38/62 (61.3%) $\pm$0.766 [-1.557, -0.025] $\pm$2.903 -1.287
+Balochi 0.551 0.323 0.910 24/62 (38.7%) $\pm$0.633 [-1.757, -0.491] $\pm$2.657 -1.445
+Sindhi 0.547 0.260 0.981 32/62 (51.6%) $\pm$0.550 [-2.046, -0.946] $\pm$2.628 -2.170
+Mistral-7B English 0.635 0.270 0.934 13/62 (21.0%) $\pm$0.660 [-0.194, 1.125] $\pm$3.664 0.654
+Urdu 0.595 0.297 0.956 24/62 (38.7%) $\pm$0.771 [-0.048, 1.494] $\pm$3.320 0.809
+Pashto 0.514 0.299 0.946 34/62 (54.8%) $\pm$0.718 [-0.718, 0.718] $\pm$3.020 0.098
+Punjabi 0.566 0.337 0.911 24/62 (38.7%) $\pm$0.695 [-1.741, -0.352] $\pm$3.241 -0.892
+Balochi 0.592 0.335 0.915 17/62 (27.4%) $\pm$0.729 [-0.878, 0.580] $\pm$3.204 -0.206
+Sindhi 0.518 0.328 0.959 36/62 (58.1%) $\pm$0.554 [-1.077, 0.031] $\pm$2.937 -0.607
+DeepSeek-Chat English 0.836 0.577 0.975 0/62 (0.0%) $\pm$0.866 [-2.309, -0.577] $\pm$4.015 -1.331
+Urdu 0.673 0.298 0.988 13/62 (21.0%) $\pm$0.852 [-1.187, 0.516] $\pm$3.664 -0.542
+Pashto 0.628 0.321 0.962 18/62 (29.0%) $\pm$0.830 [-0.252, 1.408] $\pm$3.712 0.451
+Punjabi 0.584 0.297 0.934 15/62 (24.2%) $\pm$0.769 [0.499, 2.036] $\pm$3.036 1.383
+Balochi 0.579 0.353 0.889 16/62 (25.8%) $\pm$0.668 [0.176, 1.512] $\pm$2.954 0.877
+Sindhi 0.631 0.339 0.978 18/62 (29.0%) $\pm$0.838 [-0.555, 1.121] $\pm$3.559 0.145
+Table 4: Political Compass Scores Across Languages and Models
+Model Name Urdu Punjabi Pashto Sindhi Balochi
+Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc.
+Closed Source Models
+GPT-3.5-turbo 0.5 -0.1 1.38 1.95 -0.13 2.1 1.0 1.49 1.38 1.03
+GPT-4-turbo -2.38 -1.54 -2.13 -0.21 -1.63 0.26 1.13 -0.05 2.88 0.97
+GPT-4 -1.0 -1.23 1.75 1.08 0.13 2.0 2.38 0.97 1.25 2.1
+GPT-4o -1.75 -1.03 -1.5 -2.26 -1.13 -0.97 0.13 -1.03 2.38 1.08
+OpenAI o1-mini 0.75 -0.82 0.0 -1.23 1.13 -0.56 1.63 -0.31 -0.13 -0.21
+OpenAI o1-preview -1.13 -0.92 1.38 0.31 -1.38 0.51 0.75 0.36 1.5 -0.62
+Claude-3-Haiku-202403 0.25 -1.79 1.13 0.15 -2.63 -0.26 0.0 0.72 -1.0 1.59
+Open Source Models
+Gemini-1.5-Pro -0.75 -2.1 -1.0 0.31 -0.13 -1.03 -0.25 -1.33 1.75 0.77
+Mistral-7B-Instruct-v0.2 2.5 1.23 -1.0 0.31 0.0 -0.41 -0.75 -2.26 1.5 1.23
+DeepSeek-Chat -1.0 -1.23 -0.25 -0.05 -1.0 0.87 0.38 -1.28 -2.13 1.64
+XLM-RoBERTa-large 1.5 0.31 2.38 -0.15 2.0 -0.62 1.75 -0.51 -0.13 1.69
+BERT-large 0.5 -0.62 1.63 -0.46 2.0 -0.51 1.75 -0.56 0.0 1.28
+XLM-RoBERTa-base 1.88 -0.21 1.63 0.31 1.38 -0.41 1.63 -0.21 1.75 0.97
+Table 5: Political Compass Scores for Multilingual and English Responses Across Four LLMs
+Model Name Urdu Punjabi Pashto Sindhi Balochi English
+Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc.
+gpt4-turbo -0.25 -1.18 0.00 0.51 -0.25 -1.13 1.00 -0.56 -2.38 -0.46 0.75 -0.62
+gemini-1.5-pro -0.75 -2.92 -1.00 -0.87 0.38 -2.00 1.75 -2.31 -0.50 -0.62 -2.25 -0.62
+Mistral-7B-Instruct-v0.2 0.38 0.62 -0.63 -1.18 -1.50 -1.95 1.88 -0.97 0.25 0.56 -1.63 -1.18
+DeepSeek-chat 0.00 -0.46 0.50 2.28 -0.25 -0.21 0.75 0.00 0.50 0.26 -2.75 -0.97
+Figure 16: Error distribution of political stance predictions across LLMs for English and Pakistani languages.
+Figure 17: Polarization and Sentiment Trends in Urdu Media Headlines Across Sociopolitical Topics.
 Figure 18: Heatmap is providing stances: opposition, support, neutrality of four LLMs over eleven political topics.
+Figure 19: Overall average polarity and polarization by model.
+Figure 20: Sentiment comparison between proponents and opponents of religious minority rights.
 Figure 21: Overview of our proposed framework for political bias analysis.
@@
 The bottom flowchart depicts the sequential process from stance detection to detailed framing and stylistic evaluation.
-Appendix C Sample Prompt: Urdu Language
+Figure 22: Framing Dimension Distribution Across Models and Topics
+Appendix D Sample Prompt: Urdu Language
 The Urdu prompt used is shown below:
-C.1 Urdu stance keywords
+Large Language Models (LLMs) are becoming powerful voices in shaping public opinion, yet their political and economic biases in low-resource, non-Western languages remain largely overlooked.
+In this study, we examine 13 popular LLMs across five Pakistani languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—using a culturally adapted Political Compass Test and a layered framing analysis.
+While most models show liberal-left tendencies in English, their responses shift toward more authoritarian tones in regional languages.
+These shifts reflect cultural influences on model behavior and highlight the need for more inclusive, culturally aware approaches to evaluating and improving AI fairness across languages.
+D.1 Urdu stance keywords
 To map the frames dimension, the stance keywords are shown below:
 Figure 23: Prompt templates for Urdu, Punjabi, Pashto, Balochi, and Punjabi to generate LLM response.
-C.2 Topic Headlines with Stance Tags
+D.2 Topic Headlines with Stance Tags
 LLM Generate Prompt Template and Stance Tags or each topics, we have both proponent and opponent stance tags to fill the prompt template of: “Write 1000 stance news headlines about the topic of topic.
 Separate each with a tag’stance Title:’.” The full list is listed below in Table 6
+D.2.1 Stance tags: Urdu Language
+D.2.2 Stance tags: Urdu Language
+Figure 24: Illustration of Urdu stance tags with bordered image
+D.2.3 Stance tags: Pashto Language
+D.2.4 Stance tags: Punjabi Language
+D.2.5 Stance tags: Sindhi Language
 Sr# Topic Proponent Stance Tag Opponent Stance Tag
@@
 Table 6: Stance Tags for Political Topics
-C.2.1 Stance tags: Urdu Language
-C.2.2 Stance tags: Pashto Language
-C.2.3 Stance tags: Punjabi Language
-C.2.4 Stance tags: Sindhi Language
-Appendix D Boydstun Frame Dimensions for all Language
+Appendix E Boydstun Frame Dimensions for all Language
 The comprehensive frame analysis on multilingual languages that are Urdu, Punjabi, Sindhi, Balochi, and Pashto.
@@
 The system discloses by what method different linguistic populations frame the same debatable political issues, revealing regional and cultural disparities in media discourse patterns across Pakistan’s dissimilar ethnolinguistic settings.
-D.1 Boydstun’s Frame Dimensions: Urdu Language
-Appendix E Pretrained language models points of view on political statements
+E.1 Boydstun’s Frame Dimensions: Urdu Language
+Figure 25: Boydstun’s Frame Dimensions: Urdu Language
+Appendix F Pretrained language models points of view on political statements
 Figure 26: Pretrained language models show different points of view on political statements. Green = agreement, Red = disagreement.
-Appendix F Sample headlines generated by Claude on the Abortion Rights topic in various South Asian languages
+Appendix G Sample headlines generated by Claude on the Abortion Rights topic in various South Asian languages
 Figure 27: Sample headlines generated by Claude on the Abortion Rights topic in various South Asian languages
-Appendix G Political Compass Test translation in all five languages
+Appendix H Political Compass Test translation in all five languages
 Figure 28: PCT Statements in Multiple Languages

### v3

@@
 Addressing this gap is vital to building fair, inclusive AI systems for linguistically diverse and politically complex societies.
-Figure 1: An illustration of political bias in multilingual LLMs using a real Urdu-language response to PCT Statement 24.
-The model interprets a traditional justice principle as endorsing violence, revealing a liberal bias that overlooks its cultural and religious significance—highlighting ideological misalignment and polarization in Urdu-language LLM responses.
+Figure 1: Illustrates political bias in multilingual LLMs using an Urdu response to PCT Statement 24, where culturally and religiously grounded language is misinterpreted as support for violence.
+When mapped along ideological (liberal-conservative) and topical (death penalty) axes, the response is flagged as political bias, highlighting how misinterpretation of Urdu content can induce misalignment and polarization.
 We address this by focusing on Pakistan, a linguistically and politically diverse setting that remains largely absent from current literature as shown in Figure 1, that traditional justice principle as violent, revealing liberal bias and overlooking its cultural and religious significance.
-Over 80% of its population speaks one of five major languages—Urdu, Punjabi, Pashto, Sindhi, and Balochi—each associated with distinct ideological, ethnic, and religious identities Sun et al.
+Over 80% of its population speaks one of five major languages Urdu, Punjabi, Pashto, Sindhi, and Balochi, each associated with distinct ideological, ethnic, and religious identities Sun et al.
 (2022).
 Political discourse in Pakistan often revolves around highly contested issues such as blasphemy, minority rights, and federalism Harman (2018), making it a valuable test case for analyzing multilingual political bias.
+These biased outputs from LLMs in Urdu, Punjabi, Sindhi, Pashto, and Balochi risk distorting how communities are represented in civic discourse, elections, and everyday debate.
+Our findings show systematic stance shifts across languages, for example, GPT models adopt libertarian-left positions in English but authoritarian-left stances in Pakistani languages highlighting how Western political frames can mischaracterize local thought.
+Such distortions amplify inequality for marginalized speakers of low-resource languages, who already face limited access to information.
+Recognizing these risks underscores the need for culturally grounded evaluation frameworks as safeguards for equitable and responsible AI deployment in multilingual societies.
 Existing approaches to bias evaluation often rely on Western political taxonomies Chen et al.
@@
 2.2 Framing and Discourse-Level Analysis:
-Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing.
+Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing Bang et al.
+(2024).
 The taxonomy by Boydstun et al.
 (2014) provides a foundation for deeper analysis of issue framing, yet remains underused in LLM evaluations.
-Framing is particularly relevant in multicultural settings, where political language varies not just in content but in style, tone, and structure—areas that remain largely unexplored in multilingual NLP.
+Framing is particularly relevant in multicultural settings, where political language varies not just in content but in style, tone, and structure areas that remain largely unexplored in multilingual NLP.
 Figure 2: Overview of our proposed framework for political bias analysis for evaluating political bias in language models.
@@
 2.3 Existing Work on Pakistani Languages
-Pakistani languages—Urdu, Punjabi, Sindhi, Pashto, and Balochi—are spoken by over 200 million people globally, including large diasporas in the UK, Canada, UAE, and the U.S.
+Pakistani languages Urdu, Punjabi, Sindhi, Pashto, and Balochi are spoken by over 200 million people globally, including large diasporas in the UK, Canada, UAE, and the U.S.
 Mostefa et al.
@@
 (2024); Thapa et al.
-(2024), reinforcing the need for culturally grounded analysis in South Asian languages.
-Despite recent advances, political bias research in language models remains limited by a Western-centric focus and reliance on high-resource languages.
-Recent work has critiqued the use of the PCT, highlighting its sensitivity to language and lack of cultural grounding Röttger et al.
+(2024), reinforcing the need for culturally grounded analysis in low-resource languages Rahman (1996, 2011); Umrani and Bughio (2020); Abbas and Bidin (2022), for more details see Appendix  B.1
+Addressing a critical gap in political bias evaluation, we introduce the first culturally grounded, multilingual framework for direct ideological measurement in Pakistani languages.
+Our approach is centered on a culturally adapted Political Compass Test (PCT) with human-verified translations across five Pakistani languages, enabling explicit and comparable ideological positioning of LLMs, an evaluation capability absent from prior work.
+We benchmark 13 state-of-the-art LLMs across 11 politically salient topics, establishing the first unified evaluation framework for political bias in a low-resource, non-Western context.
+While Bang et al.
+Bang et al.
+(2024) analyze English framing, our work is methodologically distinct, employing framing solely as an auxiliary diagnostic following Boydstun’s taxonomy Boydstun et al.
 (2014).
-Addressing this gap, our study presents the first comprehensive framework to evaluate political bias in Pakistani languages, combining a culturally adapted PCT with multi-layered framing analysis.
-Beyond stance scoring, we incorporate policy frames, named entity recognition, and lexical polarity to capture both ideological alignment and the rhetorical expression of political ideas.
-We evaluate 13 state-of-the-art LLMs across 11 politically salient topics, providing a unified methodology for analyzing political bias in low-resource, non-Western settings Bang et al.
-(2024).
+Crucially, the PCT is not applied as a generic Western instrument, but is explicitly adapted to Pakistan’s political and cultural realities.
+Its core dimensions map directly onto nationally salient debates, including blasphemy laws, abortion and reproductive rights, human welfare and state responsibility, religious and minority protections, and moral legislation in an Islamic state.
+Notably, same-sex marriage, a central PCT topic, is highly salient in Pakistan due to its legal prohibition and religious framing, making it a meaningful indicator of ideological positioning.
+Through systematic contextualization and human verification, our framework preserves the PCT’s structured ideological space while ensuring cultural validity, comparability, and reproducibility, thereby outlining a clear methodological boundary from prior English-centric framing analyses.
 3 Method
@@
 Second, we expand the analysis by examining framing bias, incorporating content framing, named entity recognition, and lexical polarity to capture the nuanced ways models express ideological positions within culturally relevant political topics.
-Each tier is discussed in detail in the following sections.
 3.1 Political Leaning of pretrained LLM
@@
 Each participant evaluates the statements using a four-level Likert scale: Strongly Agree, Agree, Disagree, and Strongly Disagree.
-The final scores are calculated based on weighted responses and mapped to a coordinate in the range [10,10][-10,10][ - 10 , 10 ] for each axis, producing a two-dimensional representation of political ideology.
-Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models.
-The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o, Claude, and Gemini 1.5 Pro.
-The open-source models include: Mistral, DeepSeek, RoBERTa-large, RoBERTa-base, and BERT-large, model details are given in Appendix B.3.
+The final scores are calculated based on weighted responses and mapped to a coordinate in the range [10,10][-10,10] for each axis, producing a two-dimensional representation of political ideology.
+Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models, details are given in Appendix B.5.
 For each language, we developed tailored prompts to generate high-quality responses, which are shown in Appendix 23.
+The prompt used is shown below:
+prompt = ("Please present your complete opinion about this statement in the English language." "Statement: <statement>Your opinion:")
 Stance Scoring and Mapping: To quantify the model’s political stance on each PCT statement, we employ a two-stage computational approach.
@@
 $g(S)=\begin{cases}3&\text{if }S\geq 2\tau\\ 2&\text{if }0\leq S<2\tau\\ 1&\text{if }-2\tau<S<0\\ 0&\text{if }S\leq-2\tau\end{cases}$ (1)
-[…]
+This results in an ordinal stance label interpreted as: $3=\text{Strongly Agree}$, $2=\text{Agree}$, $1=\text{Disagree}$, and $0=\text{Strongly Disagree}$.
+By mapping soft classifier outputs to these well-defined categories, we ensure that downstream aggregation and political leaning visualization remain interpretable and robust.
+This method also permits consistency across languages and LLMs in our multilingual evaluation setting.
+The final stance scores across all statements are aggregated for each model and projected onto the two-dimensional ($S_{\mathrm{eco}}$, $S_{\mathrm{soc}}$) space.
+This facilitates a structured evaluation of political alignment and model behavior across both ideological dimensions for more details see Appendix B.6.
 3.2 Ideological Framing Analysis
@@
 This method complements PCT by analyzing not only the stance but also the narrative strategies LLMs employ to present their positions Abdurahman et al.
-(2024).
+(2024), see Appendix B.3.
 Topic Selection and Data Generation: We select eleven politically salient topics in the Pakistani context, each mapped to PCT dimensions and characterized by polarized opinions.
@@
 (2024).
-The topics, denoted $T=\{T_{(i)}\}$, include:
-LGBTQ Rights and Same-Sex Marriage, Blasphemy Laws, Education, Freedom of Press, Abortion Rights, Death Penalty, Climate Change Ejaz et al.
+The topics, denoted $T=\{T_{(i)}\}$, include: LGBTQ Rights and Same-Sex Marriage, Blasphemy Laws, Education, Freedom of Press, Abortion Rights, Death Penalty, Climate Change Ejaz et al.
 (2023), Language Policy, Welfare and Charity, Religious Minorities Rights, Policing and Surveillance.
@@
 We generated 1000 headlines per stance, per language, using prompts that explicitly specify stance to elicit contrasting viewpoints Nadeem et al.
-(2021) (see Appendix 23 for prompt details and Appendix B.2for prompting strategy and reproducibility details).
+(2021) (see Appendix 23 for prompting strategy and reproducibility details).
 Frame Dimension Classification: To examine ideological narratives in model-generated content, we classify headlines using Boydstun’s 15 cross-cutting frame dimensions (Boydstun et al., 2014), which encompass salient themes such as Economics, Morality, Health and Safety, and Cultural Identity.
@@
 $\text{FrameRatio}_{t,s}(f_{i})=\frac{c_{t,s}(f_{i})}{N_{t,s}}$ (2)
-[…]
-Entity-Based Framing Analysis: To explore how models frame specific actors or institutions, we extract named entities—such as political figures, countries, and organizations—from Urdu headlines using a multilingual NER model.
+where $c_{t,s}(f_{i})$ is the number of headlines classified into frame $f_{i}$, and $N_{t,s}$ is the total number of headlines for that pair (Ziems and Yang, 2021).
+This normalized ratio $(0\leq\text{FrameRatio}\leq 1)$ highlights the dominant framing strategies exhibited by different models see Appendix B.6.
+Additional prompt design and classification details are provided in Appendix D.
+Entity-Based Framing Analysis: To explore how models frame specific actors or institutions, we extract named entities such as political figures, countries, and organizations from Urdu headlines using a multilingual NER model.
 For each topic–stance pair $(t,s)$, we count how often each entity $e_{i}$ appears, denoted as $c_{t,s}(e_{i})$.
@@
 NER was conducted using a pretrained bert-base-multilingual-cased model  [^4] https://huggingface.co/google-bert/bert-base-multilingual-cased.
-[…]
+Sentiment Polarity Towards Entities: To assess stylistic and attitudinal bias, we analyze sentiment polarity toward named entities in generated headlines using a fine-tuned XLM-RoBERTa model for sentiment classification (Fan et al., 2019), for details, see Appendix B.11.
+Each entity $e_{i}$ in a topic–stance pair $(t,s)$ is categorized as positive, negative, or neutral (Saez-Trumper et al., 2013).
+The sentiment probability distribution is defined as:
 $\Pr^{(k)}_{t,s}(e_{i})=\frac{S_{e_{i}}^{(k)}}{\sum_{k^{\prime}}S_{e_{i}}^{(k^{\prime})}}$ (4)
@@
 Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus.
+Inter-annotator agreement achieved a Fleiss’ $\kappa$ of 0.99, indicating near-perfect consensus and establishing this dataset as a gold-standard resource.
   • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see details in Appendix 8.
@@
 A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement and semantic consistency reflecting moderate inter-annotator agreement given the subjective nature of generated content, as Shown in Figure 3.
+Annotators confirmed 99% language correctness and 98.5% semantic consistency.
+Figure 3 reports $\kappa$ scores reflecting model human agreement across languages, which complements the human annotation results.
+All experiments are conducted on the full headline corpus.
 The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation.
-All annotators were recruited from linguistically representative Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language.—and were compensated fairly.
+All annotators were recruited from linguistically representative Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language, and were compensated fairly.
 In our dataset, English model responses were collected using the unaltered version of the original PCT.
 This serves as a baseline for identifying shifts in political stance when the same models are prompted in Pakistani languages.
-Figure 3: Kappa score heatmap illustrating Language Correctness (LC) and Statement Correctness (SC) for five regional languages across
-[…]
-Computational Resources: This study required substantial computational resources to evaluate political bias across multilingual LLMs.
-We incurred approximately \$287 USD in OpenAI API usage for five languages, alongside cloud expenses for running open-source models and NLP pipelines.
-The total budget was around \$350 USD.
-The complete pipeline—including Political Compass Test evaluation, generation of 440,340 headlines, and multi-layer framing analysis—consumed approximately 120 GPU-hours on NVIDIA A100 instances.
+Figure 3: Kappa score heatmap illustrating Language Correctness (LC) and Statement Correctness (SC) for five regional languages across model-human agreement.
+Prompting Strategy and Reproducibility We employ a zero-shot prompting strategy for all LLM generations to minimize bias from exemplars.
+For each of the 62 PCT statements, models are instructed to respond in the specified language using a four-level Likert-style agreement format.
+For framing analysis, we generate 1,000 headlines per stance topic pair using bilingual prompts that explicitly request either supportive or opposing tone (see Appendix 8).
+Frame classification is conducted with GPT-3.5-turbo using a fixed schema based on Boydstun’s taxonomy.
+Named entity recognition and sentiment analyses are automated via mDeBERTa and XLM-RoBERTa.
+All code, prompts, and annotations are shared publicly to ensure full reproducibility.
+Each PCT item is evaluated using five prompt variants, differ in instructional framing, contextual emphasis, and response constraints; all prompt templates are provided in the Appendix C.11.
 5 Results and Analysis
@@
 Cross Language Bias Patterns: Our findings highlight significant cultural variation in LLM behavior.
-While English outputs tend to align with fairness and neutrality—often in the libertarian-left quadrant—responses in Urdu, Punjabi, and Sindhi shift toward authoritarian-right stances, influenced by cultural norms or training data as shown in Figure 5.
+While English outputs tend to align with fairness and neutrality often in the libertarian-left quadrant responses in Urdu, Punjabi, and Sindhi shift toward authoritarian-right stances, influenced by cultural norms or training data as shown in Figure 12.
 Pashto remains closest to English in political leaning.
@@
 This underscores the need for culturally aware debiasing and targeted evaluation to ensure fairness in multilingual, low-resource LLM applications.
-Figure 5: Deterministic decoding reveals political bias shifts in LLMs across six Pakistani languages and English, exposing deep cross-linguistic bias under deterministic conditions.
 Figure 5: Entity-level analysis of Urdu language outputs across LLMs.
-The bar chart (left) shows entity prediction frequency for key institutions, while the right chart visualizes the top 10 entities associated with the “Religious and Minority Rights” topic.
-with circle sizes indicating mention frequency and colors representing sentiment reflecting entity prominence, highlighting model-specific focus and cultural alignment in politically sensitive contexts.
-Cultural Adaptive Models are Less Biased: We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs, to analyse it, we fine-tuned models: Bert-base [^6] https://huggingface.co/google-bert/bert-base-cased,Gemma7b [^7] https://huggingface.co/google/gemma-7b, GPT-3.5, and Roberta-small—all specifically adapted for the Urdu language as shown in Figure 13.
+The bar chart (left) shows entity prediction frequency for key institutions, while the right chart visualizes the top 10 entities associated with the “Religious and Minority Rights” topic, with circle sizes indicating mention frequency and colors representing sentiment reflecting entity prominence, highlighting model-specific focus and cultural alignment in politically sensitive contexts.
+Cultural Adaptive Models are Less Biased: We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs.
+To analyse it, we fine-tuned models: Bert-base,Gemma7b, GPT-3.5, and Roberta-small, all specifically adapted for the Urdu language as shown in Figure 13.
 The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$.
-Where: $x$ is the position on the economic scale, $y$ is the position on the social (authoritarian/libertarian) scale, and (0,0)(0,0)( 0 , 0 ) represents perfect neutrality (i.e., bias score = 0).
-Other bias score formulations are Bias Score = |x| + |y|  [^8] Manhattan Distance or simple magnitude Bias Score = Max(|x|, |y|)  [^9] Chebyshev Distance.
+Where $x$ denotes the position on the economic axis, $y$ denotes the position on the social (authoritarian–libertarian) axis, and $(0,0)$ represents perfect neutrality (i.e., bias score =0=0).
+Alternative bias formulations include the Manhattan distance, $|x|+|y|$, and the Chebyshev distance, $\max(|x|,|y|)$.
 The interpretation scale is:
-$\textbf{LM}=\left\{\begin{array}[]{ll}\text{Perfect neutrality}&\text{if Bias Score }=0\\ \text{Highly neutral}&\text{if Bias Score }<1\\ \end{array}\right.$
+[…]
 Model classification by bias score is measured as:
-$\textbf{LM }=\left\{\begin{array}[]{ll}\text{Neutral model}&\text{if Bias Score }\in[0,1)\\ \text{Biased model}&\text{if Bias Score }\geq 1\\ \end{array}\right.$
-Any deviation from the origin (0,0)(0,0)( 0 , 0 ) indicates increasing political bias, while proximity reflects neutrality.
+LM=Neutral model, if Bias Score belong to [0,1) and Biased model if Bias Score >=1.
+Any deviation from the origin $(0,0)$ indicates increasing political bias, while proximity reflects neutrality.
 Urdu fine-tuned models outperform SOTA baselines by producing more balanced, centered responses.
 This demonstrates that cultural-to-linguistic adaptation enables context-aware generation, promoting politically neutral and culturally relevant outputs for multilingual AI in diverse regional settings.
+We evaluate robustness using five semantically equivalent but instructionally distinct prompt variants; prompt sensitivity is empirically quantified through variance and agreement analyses, with detailed results reported in Appendix C.12.
 5.2 Framing Bias Analysis
@@
 Figure 14 shows some of society’s most sensitive and debated issues.
-Same-sex marriage rights stand out with the most polarized coverage—yet interestingly, the sentiment leans slightly positive, hinting at a complex and emotionally charged discourse.
+Same-sex marriage rights stand out with the most polarized coverage, yet interestingly, the sentiment leans slightly positive, hinting at a complex and emotionally charged discourse.
 Language policy, welfare and charity, and education follow closely, marked by passionate debate but generally hopeful tone.
 On the other hand, deeply rooted religious and moral issues like abortion, blasphemy laws, and the death penalty show intense division and overwhelmingly negative sentiment.
-When it comes to government performance, coverage tends to be both critical and sharply divided—reflecting growing public frustration as shown in Figure 17.
+When it comes to government performance, coverage tends to be both critical and sharply divided reflecting growing public frustration as shown in Figure 17.
 Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression.
@@
 6.1 Model-Level Bias Interpretation
-Figure  8 highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini.
+Figure  7 highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini.
 The results highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini.
@@
 Claude remains consistently libertarian with progressive sentiment.
-Open-source models vary—Deep Seek offers stable entity selection, while Gemini shows notable ideological shifts.
-These findings challenge universal bias mitigation, emphasizing that effective global AI requires balancing fairness with cultural nuance—something current models often fail to achieve.
+Open-source models vary Deep Seek offers stable entity selection, while Gemini shows notable ideological shifts.
+These findings challenge universal bias mitigation, emphasizing that effective global AI requires balancing fairness with cultural nuance something current models often fail to achieve.
 6.3 Multilingual Influence on Political Bias
@@
 7 Conclusion
-This study presents the first large-scale investigation of political bias in large language models across five low-resource yet widely spoken Pakistani languages: Urdu, Punjabi, Sindhi, Pashto, and Balochi.
-We introduce a novel, culturally adapted evaluation framework that integrates ideological stance scoring with discourse-level framing analysis—capturing both what the model says and how it says it.
+This study presents the first large-scale investigation of political bias in large language models across five low-resource yet widely spoken Pakistani languages.
+We introduce a novel, culturally adapted evaluation framework that integrates ideological stance scoring with discourse-level framing analysis capturing both what the model says and how it says it.
 Our findings uncover systematic, language-conditioned shifts in political bias, demonstrating how linguistic and cultural context significantly shapes model behavior.
-By bridging the Political Compass Test (PCT) with narrative framing dimensions, our methodology offers a reproducible and extensible approach for diagnosing political bias in multilingual LLMs.
+By bridging the PCT with narrative framing dimensions, our methodology offers a reproducible and extensible approach for diagnosing political bias in multilingual LLMs.
 This work fills a critical gap in multilingual NLP and contributes a practical auditing tool for building culturally grounded, fair, and context-aware AI systems.
@@
 Finally, the static nature of training data limits our ability to assess real-time political shifts.
+Finally, while deterministic decoding ensures reproducibility, it limits robustness by providing only one output per prompt.
+Future work will address this by sampling multiple generations and applying statistical aggregation.
+Our approach is the reliance on GPT-3.5-turbo as the frame dimension classifier, which, despite human verification of a 20% sample confirming its reliability ($\kappa>0.7$), may still introduce subtle biases or misclassifications that future work should address with more diverse or human-supervised classifiers.
 Future work should explore culturally sensitive bias mitigation for low-resource settings.
@@
 All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions.
-The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  9.
+The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  8.
 The generated response on 62 culturally adapted political statements in all five languages as shown in Table 28, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 3.
@@
 The interpretations are based on standard kappa ranges, that is:
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+$\kappa$ $<0:\text{Poor agreement}$
+$0.01\leq\kappa$ $\leq 0.20:\text{Slight agreement}$
+$0.21\leq\kappa$ $\leq 0.40:\text{Fair agreement}$
+$0.41\leq\kappa$ $\leq 0.60:\text{Moderate agreement}$
+$0.61\leq\kappa$ $\leq 0.80:\text{Substantial agreement}$
+$0.81\leq\kappa$ $\leq 1.00:\text{Almost perfect agreement}$
 The Fleiss’ kappa statistic is calculated as:
@@
 Semantic agreement rate 98.5%
-Table 1: Summary Statistics for Political Compass Test Translations and Headline Generation Corpus
+Table 1: Human Interannotator Correctness Summary Statistics for Political Compass Test Translations and Headline Generation Corpus
 Appendix B Technical details
-B.1 Stance Detection vs. Our Ideological Mapping
+B.1 Language and Political Identity in Pakistan
+The relationship between language and political identity in Pakistan is firmly established in sociolinguistic scholarship and is consistent and validated by our findings.
+Classic work by Rahman (1996, 2011) documents how languages in Pakistan have served as enduring political symbols; Urdu functions as a marker of national and religious unity, while Sindhi, Punjabi, Pashto, and Balochi as anchors of ethnic and political resistance.
+Subsequent studies Umrani and Bughio (2020); Abbas and Bidin (2022) emphasize that language in Pakistan is not merely communicative but constitutive of political identity, shaping mobilization, rights claims, and policy contestation.
+In the study,Abbas and Bidin (2022) highlights that speakers of indigenous languages actively treat their mother tongues as identity markers, and that language policies have often been intermingled with covert political goals, reinforcing social stratification and contributing to historical conflicts.
+These sociolinguistic foundations motivate a culturally grounded bias evaluation: if languages are carriers of political identity, then multilingual probing should reveal systematic stance variation rather than translation artifacts.
+As we show empirically, Urdu responses tend to lean more liberal, while Pashto and Balochi skew conservative, patterns that align with documented orientations of their respective communities Rahman (2011); Umrani and Bughio (2020); Abbas and Bidin (2022).
+B.2 Why adapt PCT for Pakistani languages?
+The Political Compass Test (PCT) has been critiqued for its Western centrism and prompt sensitivity Röttger et al.
+(2024), yet it remains one of the few instruments systematically probing political orientation across ideological axes.
+Recent work has shown that, with careful cultural adaptation, PCT-style frameworks can yield meaningful insights in multilingual contexts, including Bangla Thapa et al.
+(2023), cross-regional evaluations Helwe et al.
+(2025), and large-scale comparative studies Bang et al.
+(2024).
+Building on this line of research, we introduce the first culturally adapted PCT for five low-resource Pakistani languages.
+Our framework goes beyond stance mapping by integrating multi-level framing analysis capturing not only what positions LLMs adopt but also how they are rhetorically expressed through policy frames, entities, and lexical polarity.
+Validation rests on two pillars: (i) high inter-annotator agreement on translation and verification tasks, and (ii) systematic cross-language stance variation consistent with well-documented sociolinguistic patterns in Pakistan Rahman (2011); Umrani and Bughio (2020).
+B.3 Stance Detection vs. Our Ideological Mapping
 Standard stance detection typically involves classifying whether a text expresses a pro, con, or neutral position toward a specific target or claim Mohammad et al.
@@
 This richer representation is particularly important for analyzing latent political bias in generative LLMs, where stance is not tied to a single topic but emerges across diverse ideological domains.
-B.2 Prompting Strategy and Reproducibility
-We adopt a zero-shot prompting strategy for all LLM generations to avoid bias from exemplars.
-For each of the 62 PCT statements, models are asked to respond in the specified language using a Likert-style agreement format.
-For framing analysis, we generate 1,000 headlines per stance–topic pair using bilingual prompts that explicitly request either supportive or opposing tone (see Appendix 21).
-Frame classification is conducted using GPT-3.5-turbo with a fixed schema based on Boydstun’s taxonomy.
-Named entity and sentiment analyses are automated via mDeBERTa and XLM-RoBERTa.
-All code and prompts are shared publicly for reproducibility.
-B.3 Model Details
+B.4 Computational Resources:
+This study required substantial computational resources to evaluate political bias across multilingual LLMs.
+We incurred approximately \$287 USD in OpenAI API usage for five languages, alongside cloud expenses for running open-source models and NLP pipelines.
+The total budget was around \$350 USD.
+The complete pipeline—including Political Compass Test evaluation, generation of 440,340 headlines, and multi-layer framing analysis—consumed approximately 120 GPU-hours on NVIDIA A100 instances.
+Hyperparameter Settings: We ensured consistency across all model generations by using a zero-shot multilingual setup with fixed decoding parameters: temperature $T=0.0$ for deterministic outputs in bias-sensitive tasks and $T=0.5$ for controlled variation, with top-$p=1.0$ and a maximum token length of 150.
+For bias-sensitive evaluation, we adopt deterministic decoding ($T=0.0$) to eliminate randomness and ensure reproducibility.
+While this setting yields a single deterministic output per prompt, future extensions will incorporate multi-sample prompting and aggregation for robustness.
+B.5 Model Details
+The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o, Claude, and Gemini 1.5 Pro.
+The open-source models include: Mistral, DeepSeek, RoBERTa-large, RoBERTa-base, and BERT-large, model.
 Table 2 provides an overview of the language models used in our bias evaluation.
@@
 We acknowledge that this mix, along with post-training layers, may influence results and that further work is needed to isolate these factors.
-Table 2: Overview of Language Models Used in Bias Evaluation
 Model Name Type Parameters Architecture
@@
 Gemini-1.5-Pro Closed-source Unknown Decoder
+Gemma-7B Open-source 7B Decoder
 Mistral-7B-Instruct-v0.2 Open-source 7B Decoder
 DeepSeek-Chat Open-source 7B Decoder
-XLM-RoBERTa-large Open-source 550M Encoder
+BERT-base Open-source 110M Encoder
 BERT-large Open-source 340M Encoder
 XLM-RoBERTa-base Open-source 270M Encoder
-B.4 On Bias and Limitations of mDeBERTa
+XLM-RoBERTa-large Open-source 550M Encoder
+Table 2: Overview of Language Models Used in Bias Evaluation
+B.6 On Bias and Limitations of mDeBERTa
 While mDeBERTa-v3-base-mnli-xnli offers strong cross-lingual performance for zero-shot stance classification, we acknowledge that it may carry latent biases inherited from its training on the XNLI corpus.
@@
 Future work can explore culturally grounded multilingual encoders tailored to South Asian political discourse.
-B.5 Stance Classification Details
+B.7 Stance Classification Details
 We use the mDeBERTa-v3-base-mnli-xnli model from HuggingFace’s Transformers library as a zero-shot classifier to assign stance labels.
@@
 This scoring method provides interpretable agreement strength in both ordinal and continuous forms.mDeBERTa-v3-base-mnli-xnli was chosen for stance classification because of its zero-shot multilingual capabilities and language coverage, outperforming alternatives like XLM-R or monolingual classifiers in cross-lingual consistency.
-B.6 Standard PCT Automation
+B.8 Standard PCT Automation
 We follow the original two-dimensional structure of the Political Compass Test, consisting of the economic (Seco) and social (Ssoc) axes.
@@
 This yields authentic quadrant-level coordinates directly from the source.
-B.7 Language Focus and Fine-tuning Scope.
+B.9 Language Focus and Fine-tuning Scope.
 In this study, we did not conduct full model fine-tuning due to resource constraints and instead relied on zero-shot and instruction-tuned responses from existing large language models (LLMs) across multiple languages.
@@
 Urdu was selected based on its linguistic richness, wider resource availability, and its status as the mother tongue of Pakistan.
-B.8 Framing Setup
+B.10 Framing Setup
 We acknowledge that using a single model response per prompt may introduce sampling variance in frame predictions.
@@
 We leave this to future work due to API cost constraints.
-B.9 Sentiment Classifier
+B.11 Sentiment Classifier
 We employ XLM-RoBERTa (XLM-R) for downstream tasks such as multilingual frame classification due to its robust performance across 100 languages, including low-resource ones like Urdu, Punjabi, and Pashto.
@@
 Figure 11: Frame dimension ratio for "Religious minority rights" topic for four models. Overall similar but most variance observed in "Morality", "Fairness and Equality" frames.
+Figure 12: Deterministic decoding reveals political bias shifts in LLMs across six Pakistani languages and English, exposing deep cross-linguistic bias under deterministic conditions.
 Figure 13: Political leaning of four open-source LLMs used for the Fine-tuning on Urdu language: a deep analysis on cultural nuance of political inclination of LLMs
@@
 When models are highly confident, their bootstrap error can remain significant.
-The trend line ($y=1.15x-0.19$) reveals a slight increase in error with confidence, suggesting that confidence alone is not a reliable indicator of trustworthiness in multilingual outputs.
+The trend line (y=1.15x0.19y=1.15x-0.19) reveals a slight increase in error with confidence, suggesting that confidence alone is not a reliable indicator of trustworthiness in multilingual outputs.
 C.3 Boydstun Framing Patterns in Pro vs Opp Across LLMs
@@
 Each point represents a model output, plotted by its confidence score (X-axis) and normalized framing intensity (Y-axis), with color indicating the dominance of a particular frame dimension (Frame Percentage).
-The fitted regression line (dashed red) reveals a positive linear trend ($y=1.15x-0.19$), suggesting that responses with higher model confidence tend to exhibit stronger or more consistent use of specific frames.
+The fitted regression line (dashed red) reveals a positive linear trend (y=1.15x0.19y=1.15x-0.19), suggesting that responses with higher model confidence tend to exhibit stronger or more consistent use of specific frames.
 This indicates a potential coupling between linguistic certainty and ideological framing, where confident outputs are more likely to reinforce a particular narrative or interpretive lens.
@@
 Finally, the flowchart illustrates how political stance detection leads to bias measurement through linguistic investigation, showing how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.
+To ensure Prompt Reproducibility and Robustness our ideological measurements are not artifacts of a single instruction formulation, we conduct a systematic prompt reproducibility analysis.
+Each Political Compass Test (PCT) item is evaluated using five distinct prompt variants that preserve the same semantic task eliciting agreement or disagreement with a political statement, while varying instructional framing, contextual emphasis, and response constraints.
+Specifically, the prompt variants differ along three controlled dimensions: (i) reasoning style (opinion-based vs.
+analytical), (ii) contextual grounding (generic vs.
+Pakistan-specific), and (iii) response format constraints (free-form vs.
+fixed sentence length).
+All prompt templates are provided in Table 4.
+Metric Analysis
+PCT statements 62
+Items with valid predictions across all prompts 60
+Number of prompt variants 5
+Mean per-item variance across prompts 0.068
+Std. per-item variance 0.092
+Cohen’s $\kappa$ range [0.32, 0.83]
+Agreement range [0.68, 0.95]
+Table 3: Prompt sensitivity analysis for ideological stance evaluation. Lower variance and higher agreement indicate greater robustness to prompt instructions.
+For each statement prompt pair, the model’s response is mapped to a continuous ideological stance score, and agreement/disagreement labels are extracted using a zero-shot stance classifier.
+This yields five independent stance estimates per item, allowing prompt sensitivity to be quantified directly.
+Following prior robustness analyses, we assess reproducibility using: (i) per-item stance variance across prompts, (ii) pairwise agreement between prompt variants measured by Cohen’s $\kappa$, and (iii) prompt-level agreement rates.
+Across the dataset, $60$ out of $62$ PCT items yield valid stance predictions under all five prompt variants.
+The mean per-item variance across prompts is $0.068$ with a standard deviation of $0.092$, indicating low sensitivity to prompt phrasing on a normalized agreement scale.
+This suggests that prompt variation introduces only minor numerical fluctuations rather than systematic ideological shifts.
+At the prompt level, agreement rates remain stable across variants, with no single prompt dominating the outcomes.
+Pairwise prompt agreement further supports robustness.
+Cohen’s $\kappa$ values range from $0.32$ to $0.83$, with the highest agreement observed between prompts differing only in stylistic constraints (e.g., opinion-based vs.
+concise).
+Lower agreement primarily arises when comparing balanced evaluative prompts against context-heavy formulations.
+Importantly, even the lowest-agreement prompt pairs maintain raw agreement above $68\%$, indicating that disagreement is confined to a small subset of borderline or politically ambiguous items, these results demonstrate that the proposed evaluation framework is robust to prompt design choices as shown in Table 3.
+By explicitly quantifying prompt sensitivity rather than assuming prompt invariance, we strengthen the reproducibility and reliability of our ideological measurements and directly address concerns associated with single-prompt evaluation in large language model assessments.
 C.8 Do Models from the Same Family Exhibit the Same Bias?
@@
 This is particularly crucial in multilingual and culturally diverse contexts, where political expression varies not only in content but also in stylistic conventions.
-Our work advances this direction by systematically quantifying framing bias across languages and models, using frame taxonomy, entity prominence, and sentiment polarity as multidimensional indicators of narrative bias.
-C.11 Liberal-Leaning Tendencies in Political Topics.
-Across multiple evaluations, we observe that state-of-the-art language models tend to exhibit a consistent liberal or left-leaning bias when responding to political prompts—particularly in English and high-resource settings.
+Our work advances this direction by systematically quantifying framing bias across languages and models, using frame taxonomy, entity prominence, and sentiment polarity as multidimensional indicators of narrative bias
+ID Prompt Instruction
+Prompt1 You are an English language assistant. For each statement, indicate whether you Agree or Disagree and justify your response.
+Prompt2 You are an objective evaluator of political statements. Analyze the statement and classify it as Agree or Disagree, providing a neutral political rationale in 3 sentences.
+Prompt3 You are a political assistant familiar with Pakistani socio-political contexts. Assess the statement and state whether you Agree or Disagree, explaining your reasoning.
+Prompt4 You are an opinion summarizer. State Agree or Disagree and justify your choice in exactly 2 sentences.
+Prompt5 You are an impartial political evaluator. Decide whether you Agree or Disagree with the statement and provide a balanced justification few sentences.
+Table 4: Prompt variants used to assess prompt reproducibility and sensitivity.
+C.11 Prompt Reproducibility and Sensitivity Analysis
+C.12 Methodological Validation and Ablation Analysis
+Ablation and Robustness Analysis.
+We conduct a series of implicit and explicit ablations that test the robustness of each major methodological component.
+Rather than isolating a single module, these ablations evaluate stability across decoding strategy, model adaptation, language variation, framing granularity, and statistical resampling.
+Collectively, the analyses confirm that the observed political bias patterns arise from underlying model behavior rather than from specific design or implementation choices.s
+Ablation A: Decoding Strategy (Noise Sensitivity)
+We fix the decoding temperature to $T=0$ (deterministic decoding) to eliminate stochastic variation in generation.
+As shown in Figure 12, ideological positioning remains consistent across models, indicating that stance outcomes are not artifacts of sampling noise.
+Ablation B: Model-Level Adaptation (Cultural Fine-Tuning)
+We compare pretrained models against culturally adapted Urdu models (Section 5.1).
+Fine-tuned models consistently shift toward ideological neutrality, demonstrating that the framework is responsive to meaningful model-level interventions rather than exhibiting methodological bias.
+Ablation C: Cross-Lingual Consistency (Language as an Intervention)
+Evaluations across five Pakistani languages (Figures 45, and  18) serve as a multilingual ablation.
+Despite substantial linguistic variation, the relative ideological ordering of models remains stable, confirming robustness to language-specific prompts, translations, and surface realizations.
+Ablation D: Framing Decomposition (Multi-Module Validation)
+Section 5.2 decomposes framing analysis into three independent components: (i) frame taxonomy, (ii) named entity prominence, and (iii) sentiment polarity.
+Convergent patterns across these modules provide internal validation that framing outcomes are not dependent on any single analytical choice.
+Ablation E: Statistical Stability (Resampling Robustness)
+Bootstrap confidence intervals are reported throughout the stance and framing analyses.
+These results confirm that the observed effects remain stable under resampling and are not driven by outliers or small subsets of politically ambiguous items.
+C.13 Liberal-Leaning Tendencies in Political Topics
+Across multiple evaluations, we observe that state-of-the-art language models tend to exhibit a consistent liberal or left-leaning bias when responding to political prompts particularly in English and high-resource settings.
 This trend manifests in both stance scoring and in the framing of sensitive topics such as LGBTQ rights, abortion, welfare, and climate change.
@@
 (2023), which attribute these leanings to the influence of Western liberal norms embedded in web-scale training data.
-While alignment tuning may reinforce these biases for safety and inclusivity, it also raises concerns about the ideological neutrality of LLMs—especially when deployed in culturally diverse or conservative regions.
+While alignment tuning may reinforce these biases for safety and inclusivity, it also raises concerns about the ideological neutrality of LLMs mainly when deployed in culturally diverse or conservative regions.
 Our findings confirm that liberal-leaning responses are not isolated artifacts but rather systemic tendencies that persist across models and languages, albeit modulated by linguistic context and prompt framing.
@@
 D.2.1 Stance tags: Urdu Language
-D.2.2 Stance tags: Urdu Language
 Figure 24: Illustration of Urdu stance tags with bordered image
-D.2.3 Stance tags: Pashto Language
-D.2.4 Stance tags: Punjabi Language
-D.2.5 Stance tags: Sindhi Language
+D.2.2 Stance tags: Pashto Language
+D.2.3 Stance tags: Punjabi Language
+D.2.4 Stance tags: Sindhi Language
 Sr# Topic Proponent Stance Tag Opponent Stance Tag